### v1

# FALCON: Fairness Learning via Contrastive Attention Approach to Continual Semantic Scene Understanding in Open World

Abstract
Continual Learning in semantic scene segmentation aims to continually learn new unseen classes in dynamic environments while maintaining previously learned knowledge. Prior studies focused on modeling the catastrophic forgetting and background shift challenges in continual learning. However, fairness, another major challenge that causes unfair predictions leading to low performance among major and minor classes, still needs to be well addressed. In addition, prior methods have yet to model the unknown classes well, thus resulting in producing non-discriminative features among unknown classes. This paper presents a novel Fairness Learning via Contrastive Attention Approach to continual learning in semantic scene understanding. In particular, we first introduce a new Fairness Contrastive Clustering loss to address the problems of catastrophic forgetting and fairness. Then, we propose an attention-based visual grammar approach to effectively model the background shift problem and unknown classes, producing better feature representations for different unknown classes. Through our experiments, our proposed approach achieves State-of-the-Art (SOTA) performance on different continual learning settings of three standard benchmarks, i.e., ADE20K, Cityscapes, and Pascal VOC. It promotes the fairness of the continual semantic segmentation model.
1 Introduction
Figure 1: Our Fairness Learning via Contrastive Attention to Continual Semantic Segmentation. The Fairness Contrastive Clustering Loss promotes the fairness of the model while the Attention-based Visual Grammar models the unknown classes.
The semantic segmentation networks, e.g., Transformers [56] and Convolutional Neural Networks [7], learned from data with a closed-set of known classes have shown outstanding performance. However, they often suffer performance degradation when encountering novel objects or classes in new dynamic environments [19, 3, 52]. To improve their performance, several transfer learning and domain adaptation methods [26, 1, 52] were introduced to adapt trained models into deployed environments. While the former often aims to fine-tune the model on labeled data collected in the new environments, the latter adapts the model to the new domains in an unsupervised manner [1, 26]. However, these methods cannot handle novel objects well due to their close-set learning. In practice, the semantic segmentation models should be able to adaptively and continually learn the new knowledge of novel classes. It motivates the development of Continual Learning paradigm [3, 19, 40], a.k.a, Continual Semantic Segmentation (CSS), where the segmentation models are learned sequentially to new contents of data.
Far apart from prior segmentation methods [7, 56] that learn one time on static, closed-set data, Continual Learning requires the segmentation models to learn from dynamic, open-set data [19, 52]. In addition, in particular scenarios, accessing previous learning data is restricted due to privacy concerns. In CSS, three challenges have been identified, including (1) Catastrophic Forgetting, (2) Background Shift, and (3) Fairness. While the catastrophic forgetting problem [44, 23, 50] depicts the segmentation model tends to forget its knowledge when learning new data, background shift indicates the problem of classes of previous or future data (unknown classes) have collapsed into a background class [3, 19, 60]. Prior methods [19, 40, 5, 52] addressed these two problems by introducing knowledge distillation and pseudo labels. However, these methods can not handle unknown classes since they either consider these unknown classes as a background class or assign unknown pixels by a pseudo label of prior known classes [19, 52]. More importantly, the last problem, fairness, is a significant challenge that limits the performance of CSS models.
As shown in Fig. 2, the number of pixels of each class in training data have been imbalanced among classes and significantly decreased after each task. Thus, this bias influences the learning procedure and model predictions that later cause unfair predictions among classes. However, limited studies are taking the fairness problem into account. [51] presented a similar problem in domain adaptation and extended it to continual learning [52]. These methods rely on the assumption of ideal fair or balanced data distributions. However, it is not applicable in practice since the size, i.e., the number of pixels, of several classes can never be more significant than others. For example, the size of the bottle should not be more significant than the size of a car. Meanwhile, the current knowledge distillation methods [19, 39, 4] in CSS are unable to handle the fairness problem since they focus on modeling catastrophic forgetting and background shift problems. Therefore, it is essential to develop a new CSS approach to address these limitations.
Contributions of This Work: This work presents a novel Fairness Learning via Contrastive Attention Approach (FALCON) to Continual Semantic Segmentation (as shown in Fig. 1). First, we introduce a novel Contrastive Clustering Paradigm approach to Continual Learning that models the catastrophic forgetting problem. Second, by analyzing the limitation of vanilla Contrastive Clustering in biased data, we introduce a novel Fairness Contrastive Clustering loss to model the fairness problem in continual learning efficiently. Third, to effectively model the background shift problem, we introduce a new Attention-based Visual Grammar that model the topological structures of feature distribution to handle the unknown classes effectively. Finally, the ablation studies illustrate the effectiveness of the proposed approach in different aspects of fairness promotion in CSS models. Compared with prior methods, our approach achieves state-of-the-art performance on different settings of three standard benchmarks of CSS, including ADE20K, Pascal VOC, and Cityscapes.
Figure 2: The Data Class Distribution of ADE20K. The major classes occupy more than 75% of the total pixels of the dataset.
2 Related Work
Continual Semantic Segmentation Several studies have been introduced to address catastrophic forgetting and background shift problems in CSS [20, 28, 32, 18, 19, 38, 37, 34]. The common approach of CSS adopts knowledge distillation [19] and pseudo labels [3] to model catastrophic forgetting and background shift, respectively. Later, it is further improved by decoupling knowledge representations [60], modeling the inter- and intra-class knowledge [40], distinguishing the feature representations of the future classes [5, 39], reducing background confusion [59], or modeling distillation loss via the geodesic flow [48]. Another approach [4] adopted the mask-based segmentation networks [12, 13] to improve the performance of CSS models. Recent studies have introduced CSS under the unsupervised domain adaptation settings [54, 45, 46]. However, prior studies have yet well modeled the unknown classes in the open-world environments. Particularly, as previous methods [19, 4] use the pseudo labels to model the unknown classes, the future classes will be treated as a background class. Then, Joseph et al. [27] improves the unknown class modeling by using clustering but this method considers all different unknown classes as a single cluster, leading to the non-discriminative features among unknown classes.
Contrastive Learning is a common learning approach [36, 10, 9] to structure the deep feature representations in the deep latent space. Oorde et al. [36] first introduced the Noise-Contrastive Estimation (InfoNCE) learning framework. Then, Chen et al. [9] presented SimCLR, a self-supervised contrastive learning approach to improve the representation power of Residual Networks. He [25] proposed a Momentum Contrast framework for unsupervised representation learning. Later, it was further improved by using MLP projection head [10] and extended to improve the self-supervised training process of vision transformers [11]. Cui et al. [16] introduced a supervised parametric contrastive learning loss to address the long-tailed recognition. Li et al. [30] adopted contrastive learning to develop the one-stage online contrastive clustering method. Radford et al. [41] presents a contrastive framework to learn the vision-language model. Later, several methods also adopted this framework to vision-language pretraining [41, 29].
Imbalanced and Fairness Learning The early methods utilized the balanced Softmax loss [43] to alleviate the impact of imbalanced data distribution. Later, Wang et al. [55] introduced a Seesaw loss to re-balance the contributions of positive and negative instances via the mitigation and compensation modules. Ziwei et al. [31] introduced a dynamic meta-embedding to model the imbalanced classification problem. Chu et al. [14] reduce the bias in the segmentation model by presenting a new stochastic training scheme. Szabo et al. [49] presented a tilted cross-entropy loss to promote class-relevant fairness. However, there are limited studies that address the fairness problem in CSS. Truong et al. [51] introduced a fairness domain adaptation approach to semantic segmentation and later extended it into continual learning setting [52]. However, these methods [51, 52] rely on the assumption of ideal balanced data which could not be achieved by nature. To address the limitations in prior work, this paper will introduce a novel approach to effectively model the fairness problem and unknown classes in the continual learning setting.
Figure 3: The Proposed Open-World Fairness Continual Learning Framework.
3 The Proposed FALCON Approach
CSS aims to learn a segmentation network $F$ on sequence data $\mathcal{D}=\{\mathcal{D}^{1},...,\mathcal{D}^{T}\}$ where $T$ is the number of learning steps. At learning step $t$, the model $F$ encounters a dataset $\mathcal{D}_{t}=\left\{(\mathbf{x}^{t},\mathbf{\hat{y}}^{t})\right\}$ where $\mathbf{x}^{t}\in\mathbb{R}^{H\times W\times 3}$ is the image and $\mathbf{y}\in\mathbb{R}^{H\times W}$ is a segmentation label of $\mathbf{x}^{t}$. The ground truths at learning step $t$ only consist of current classes $\mathcal{C}^{t}$, while the class labels of the previous $\mathcal{C}^{1...t-1}$ and future steps $\mathcal{C}^{t+1...T}$ are collapsed into a background class. Formally, learning the CSS model at step $t$ can be formed as Eqn. (1).
$\footnotesize\theta_{t}^{*}=\arg\min_{\theta_{t}}\mathbb{E}_{\mathbf{x}^{t},\mathbf{\hat{y}}^{t}\in\mathcal{D}^{t}}\left[\mathcal{L}_{CE}\left(\mathbf{y}^{t},\mathbf{\hat{y}}^{t}\right)+\lambda_{CL}\mathcal{L}_{CL}\left(F(\mathbf{x}^{t})\right)\right]$ (1)
where, $\mathbf{y}^{t}=F(\mathbf{x}^{t},\theta_{t})$, $\theta_{t}$ is the parameter of $F$ at current learning step $t$, $\mathcal{L}_{CE}$ is the cross-entropy loss, $\lambda_{CL}$ is the balanced weight. and $\mathcal{L}_{CL}$ is the CSS objective. At learning step $t$, the segmentation model $F$ is required to be able to predict both previously learned classes $\mathcal{C}^{1...t-1}$ and current new classes $\mathcal{C}^{t}$. Under this learning scenario, three challenges have been identified, i.e., Catastrophic Forgetting, Background Shift, and Fairness. Several prior methods were presented to model the two first issues in CSS using knowledge distillation [19, 4]. The last issue has not been well addressed yet due to its challenges [52]. Prior methods [47, 3, 19, 60, 4] adopt knowledge distillation to design $\mathcal{L}_{CL}$. However, this method prevents the CSS model from diverging knowledge learned previously, therefore resulting in limiting the ability to adopt new knowledge [52]. In addition, these methods have not addressed fairness and background shift problems due to their dedicated design for maintaining knowledge via distillation [19, 4, 39]. Therefore, to address these problems, we introduce a novel Fairness Learning via Contrastive Attention Approach to CSS.
3.1 Continual Learning via Contrastive Clustering
Apart from prior methods [19, 4, 39], our CSS is defined as Contrastive Clustering Learning. Given a set of centroid vectors $\{\mathbf{c}_{i}\}_{i=1}^{N_{K}+N_{U}}$ where $N_{K}=|\mathcal{C}^{1..t}|$ and $N_{U}$ is the number of known and unknown classes up to current learning tasks. Prior work [27, 52, 4] often defined the number of unknown classes as $1$ where background classes are considered as a single unknown class. Formally, our Contrastive Clustering Learning for CSS can be defined as Eqn. (2).
$\footnotesize\begin{split}\mathcal{L}_{CL}\left(F(\mathbf{x}^{t})\right)&=\sum_{\mathbf{c}_{i}}\mathcal{L}_{Cont}(\mathbf{F}^{t},\mathbf{c}_{i})\\ &=\sum_{\mathbf{c}_{i}}\sum_{h,w}-\phi(\mathbf{f}^{t}_{h,w},\mathbf{c}_{i})\log\frac{\exp(\mathbf{f}^{t}_{h,w}\times\mathbf{c}_{i})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c}_{i})}\end{split}$ (2)
where $\mathbf{F}^{t}\in\mathbb{R}^{H\times W\times D}$ is the feature maps extracted from the input image $\mathbf{x}^{t}$ by the segmentation network $F$, $\mathbf{f}^{t}_{h,w}\in\mathbb{R}^{D}$ is the feature at the pixel location $(h,w)$ of features $\mathbf{F}^{t}$, $\sum_{\mathbf{f}^{\prime}}$ means the summation over all feature representations $\mathbf{f}^{\prime}\in\mathbb{R}^{D}$, and $\phi:\mathbb{R}^{D}\times\mathbb{R}^{D}\to[0,1]$ is the function that determines either $\mathbf{f}^{t}_{h,w}$ belongs to the cluster $\mathbf{c}_{i}$ or not.
By defining CSS as contrastive clustering learning, the knowledge of the segmentation model has been well maintained via the cluster vectors $\mathbf{c}$. Then, minimizing Eqn. (2) will separate the representations of different classes while gathering the features of the same class into the same cluster. As the cluster vectors $\mathbf{c}$ of the old classes $\mathcal{C}^{1..t-1}$ have been well learned to represent their knowledge, these vectors are frozen at learning step $t$ to maintain the knowledge representations of previous classes to address the catastrophic forgetting problem. To effectively learn cluster vectors $\mathbf{c}$, the cluster vector $\mathbf{c}$ will periodically updated after each $M$ steps by the momentum update [27, 52, 25] based on the features $\mathbf{f}^{t}_{h,w}$ assigned to cluster $\mathbf{c}$. However, there are two major problems in contrastive clustering learning. First, since the training data in CSS suffer the bias among classes as shown in Fig. 2, this bias will influence Eqn. (2) and cause the unfair predictions. Second, as the function $\phi$ requires the labels to determine the features belonging to clusters, it limits the ability to model the unknown classes where the labels are not available. Therefore, Secs. 3.2-3.3 will present a novel approach to tackle these problems.
3.2 Fairness Contrastive Clustering Learning
While the contrastive clustering learning defined in Eqn. (2) promotes the compact representations of features around their clusters, inspired by [62, 16, 8], we observe that the imbalanced class distribution will influence unfair behaviors among classes. In particular, for simplicity, we consider $\{\mathbf{f}^{t}_{i}\}_{i=1}^{L}$ is the set of features that belong to the cluster $\mathbf{c}$ at learning step $t$ (i.e., $\phi(\mathbf{f}_{i}^{t},\mathbf{c})=1$) and $L$ is the number of features (in this case, $L$ is the total number of pixels belong to the class of cluster $\mathbf{c}$). Let us define the enforcement between the feature $\mathbf{f}^{t}_{t}$ and the cluster $\mathbf{c}$ as $\ell_{i}=\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}$. Hence, the lower the value of the enforcement $\ell_{i}$ is, the more compact the representation of visual features and clusters is. Then, the contrastive clustering learning loss in Eqn. (2) of the entire cluster $\mathbf{c}$ can be defined as Eqn. (3).
$\small\mathcal{L}_{Cont}(;,\mathbf{c})=-\sum_{i=1}^{L}\log\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}=-\sum_{i=1}^{L}\log\ell_{i}$ (3)
Proposition 1: If the contrastive clustering loss $\mathcal{L}_{Cont}(;,\mathbf{c})$ achieves the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converge to $\ell_{i}=L^{-1}$.
Proposition 1 has implied that the class with more samples will result in a lower value of the enforcement and produce a more compact representation while the class having fewer samples will be more scattered in the feature space due to the higher value of the enforcement. In particular, let $L_{major}$ and $L_{minor}$ be the number of samples of the major and minor class where $L_{major}>>L_{minor}$. Then, based on Proposition 1, the enforcement between features and the cluster of the major class will be significantly lower than the one of the minor class, i.e., $L_{major}^{-1}<<L_{minor}^{-1}$. Therefore, a direct adoption of the contrastive clustering loss in Eqn. (2) will result in an unfair CSS model. In addition, for classes in the minority group, the weak enforcement results in the feature presentations of classes being far away from their clusters. Thus, the model will produce non-discriminative features compared to the ones in the majority group. Moreover, if the loss is applied to the cases of unknown labels, these feature representations can be scattered in the latent space and pulled into the incorrect clusters due to weak enforcement between features and clusters (Fig. 4).
To address the unfair problem in contrastive clustering learning, inspired by [62, 16, 8], we introduce a scaling factor $\alpha$ and a learnable transition vector $\mathbf{v}$ for each cluster $\mathbf{c}$ (all clusters have the same value of $\alpha$ but different vector $\mathbf{v}$). Our Fairness Contrastive Clustering Learning Loss for the entire cluster in Eqn. (3) can be re-formed as:
$\footnotesize\mathcal{L}^{\alpha}_{Cont}(;,\mathbf{c})=-\alpha\sum_{i=1}^{L}\log\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}-\log\frac{\exp(\mathbf{v}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}$ (4)
Intuitively, the scaling factor $\alpha$ will help to re-scale the impact of the enforcement in learning, and the transitive vector $\mathbf{v}$ assists in translating the center cluster into the proper position of the latent space. This action promotes the compactness of clusters in the minority group.
Proposition 2: If the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}(;,\mathbf{c})$ achieves the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converge to $\ell_{i}=(\alpha^{-1}+L)^{-1}$. Proofs of Propositions 1-2 are provided in the supplementary.
Under the Proposition 2, when the value of $\alpha$ is small, the divergence of the enforcement between major and minor classes will be smaller, i.e., $||(\alpha^{-1}+L_{major})^{-1}-(\alpha^{-1}+L_{minor})^{-1}||<||L^{-1}_{major}-L^{-1}_{minor}||$. Fig. 4 has illustrated the impact of fairness contrastive clustering loss. Therefore, our designed proposed fairness contrastive loss has effectively addressed the fairness issue in Eqn. (2). It should be noted that although the smaller $\alpha$ results in the fairer enforcement varied from major to minor classes. However, if the value of scaling factor $\alpha$ is too small, the contrastive clustering loss will rely more on the enforcement of the transitive vector $\mathbf{v}$, and the distribution of features $\mathbf{f}^{t}_{i}$ around its cluster $\mathbf{c}$ will be scattered due the weak enforcement caused by small $\alpha$. Therefore, the value of scaling factor $\alpha$ in practice should be carefully selected.
3.3 Open-world Unknown Class Modeling
An ideal CSS approach must be able to model the unknown classes without supervision, especially in open-world environments [27, 52]. Prior studies have adopted the pseudo-label strategies [19, 4] based on the model predictions to assign labels for seen classes while unseen classes have been ignored, thus resulting in non-discriminative features. [27, 52] improved the background modeling by using an additional prototypical representation for unknown classes. However, these approaches consider different unknown classes as one (i.e., $N_{U}=1$) resulting in non-distinguished representations of different unknown classes. Thus, modeling function $\phi$ in Eqn. (2) without supervision of different unknown classes (i.e., $N_{U}>1$) is challenging.
Figure 4: The Enforcement Loss of Vanilla Contrastive Clustering $\mathcal{L}_{Cont}$ and Fairness Contrastive Clustering $\mathcal{L}^{\alpha}_{Cont}$ on Pascal VOC. Since $\mathcal{L}_{Cont}$ suffers severe biased between major and minor classes, its clusters of minor classes remain scattered. Meanwhile, our $\mathcal{L}^{\alpha}_{Cont}$ produces a more uniform loss among classes, thus promoting the fairness and compactness of clusters.
Although modeling $\phi$ to determine the single feature $\mathbf{f}$ belonging to the cluster $\mathbf{c}$ is challenging, prior studies in clustering [35, 58, 57] have suggested that determine a set of features $\{\mathbf{f}^{t}_{i}\}_{i=1}^{M}$ belonging to cluster $\mathbf{c}$ should be easier. This derives from the fact that even though the feature representations of different classes are different, the distributions of features around its cluster (termed as Visual Grammar) in the feature space should be similar among classes or clusters. As a result, by learning the distribution of features and their clusters, the model $\phi$ can determine whether a feature belongs to a cluster. Then, by learning the model $\phi$ on prior known clusters and features, the knowledge of $\phi$ can be adaptively applied to unknown clusters. Fig. 5 illustrates our visual grammar model of the cluster distributions.
Limitations of Prior Clustering Methods The traditional methods in clustering, e.g., KNN or density-based clustering [21], remain limited to noisy features leading to producing the incorrect cluster assignment. Meanwhile, the modern clustering methods, e.g., Graph Neural Networks (GNNs) [58, 57], require a large memory to build the affinity graph for clusters. In addition, GNNs often learn the local structures of graphs (or clusters) and accumulate them via the aggregation layers. Hence, the global structures of the clusters, i.e., visual grammar, are not well modeled by GNNs [35]. Therefore, to address these limitations, we introduced a new Attention-based Visual Grammar approach to efficiently model the distribution of features and their clusters via the self-attention mechanism [53].
Remark 1: Given a center $\mathbf{c}$ and a set of $M$ features $\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}$ where $\mathbf{f}^{\mathbf{c}}_{i}$ denotes the feature $\mathbf{f}_{i}$ belonging to the cluster $\mathbf{c}$, and $\forall i\in[1..M-1]:\cos(\mathbf{f}_{i}^{\mathbf{c}},\mathbf{c})\geq\cos(\mathbf{f}_{i+1}^{\mathbf{c}},\mathbf{c})$ the Visual Grammar of the cluster $\mathbf{c}$ parameterized by $\Theta$ can be defined as Eqn. (5).
$\footnotesize\begin{split}&\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{f}^{\mathbf{c}}_{1},\mathbf{f}^{\mathbf{c}}_{2},...,\mathbf{f}^{\mathbf{c}}_{M},\mathbf{c},\Theta)\right]\\ =&\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{\Delta}^{\mathbf{c}}_{1},\mathbf{\Delta}^{\mathbf{c}}_{2},...,\mathbf{\Delta}^{\mathbf{c}}_{M},\mathbf{c},\Theta)\right]\end{split}$ (5)
where $\Delta^{\mathbf{c}}_{i}=\mathbf{f}^{\mathbf{c}}_{i}-\mathbf{c}$. Eqn. (5) defines the visual grammar of the cluster by modeling the feature distribution of $\mathbf{f}_{i}^{\mathbf{c}}$ and its cluster center $\mathbf{c}$. Let $\phi:\mathbb{R}^{(M+1)\times D}\to[0,1]^{M}$ be a function receiving a center $\mathbf{c}$ and a set of $M$ features $\{\mathbf{f}_{i}\}_{i=1}^{M}$ $\left(\cos(\mathbf{f}_{i},\mathbf{c})\geq\cos(\mathbf{f}_{i+1},\mathbf{c})\right)$ to determine whether $\mathbf{f}_{i}$ belonging to $\mathbf{c}$, i.e., $\mathbf{u}=\phi(\mathbf{\Delta}_{1},\mathbf{\Delta}_{2},...,\mathbf{\Delta}_{M},\mathbf{c})$ where $\mathbf{\Delta}_{i}=\mathbf{f}_{i}-\mathbf{c}$, $\mathbf{u}=[u_{1},u_{2},...,u_{M}]$ and $u_{i}=1$ denotes $\mathbf{f}_{i}$ belong to cluster $\mathbf{c}$ and vice versa. Hence, the visual grammar model in Eqn. (5) can be modeled by the network $\phi$ as follows with parameter $\Theta$ as follows:
$\footnotesize\begin{split}\Theta^{*}&=\arg\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{u}|\mathbf{\Delta}_{1},\mathbf{\Delta}_{2},...,\mathbf{\Delta}_{M},\mathbf{c},\Theta)\right]\end{split}$ (6)
Eqn. (6) aims to model the distribution of features around its cluster by learning the correlation of relatively topological structures $\mathbf{\Delta}_{i}$ of features $\mathbf{f}_{i}$ around cluster $\mathbf{c}$. Then, based on knowledge of the cluster distribution, the model $\phi$ is able to determine whether a feature $\mathbf{f}_{i}$ belongs to cluster $\mathbf{c}$. Hence, it is essential that the model $\phi$ has the ability to exploit the correlation between features $\mathbf{f}_{i}$ and cluster $\mathbf{c}$ to learn the topological structure of visual grammar. Therefore, we adopt the self-attention mechanism [53, 35] to efficiently model these feature correlations. Particularly, the model $\phi$ is formed by $L_{\phi}$ blocks of self-attention as follows:
$\footnotesize\begin{split}\mathbf{z}_{0}&=\operatorname{LN([\mathbf{\Delta}_{1},...,\mathbf{\Delta}_{M},\mathbf{c}])}+\boldsymbol{\beta},\quad\mathbf{a}_{l}=\mathbf{z}_{l}+\operatorname{MHSA}(\mathbf{z}))\\ \mathbf{z}_{l+1}&=\mathbf{a}_{l}+\operatorname{MLP}(\operatorname{LN}(\mathbf{a}_{l})),\quad\quad\quad\quad\mathbf{u}=\operatorname{Proj}(\mathbf{z}_{L_{\phi}})\end{split}$ (7)
where $\boldsymbol{\beta}$ is the positional embedding, $\operatorname{LN}$ is Layer Normalization, $\operatorname{MHSA}$ is multi-head self-attention, $\operatorname{MLP}$ is the multi-layer perception, and $\operatorname{Proj}$ is the linear projection. By using Transformers, the correlation of cluster distributions can be well modeled by the self-attention mechanism.
Figure 5: The Proposed Visual Grammar Model.
Cluster Assignment via Visual Grammar Instead of assigning the clusters based on the model prediction [19, 4, 39] or nearest cluster [52, 27] that are less effective, the cluster assignment in our approach will be performed by the visual grammar model, i.e., the visual grammar model will consider the $M$ closest features around cluster $\mathbf{c}$ to assign the cluster for these features. Then, the cluster assignments are used to compute our Fairness Contrastive Clustering loss. In addition, following common practices [19, 4, 39], we improve background shift modeling by using the cluster assignments of features as the pseudo labels of pixels.
Unknown Cluster Initialization Prior work [52, 27] initialized a single unknown cluster ($N_{U}=1$), thus resulting in producing non-discriminative class-wise features. However, there should be more than a single unknown cluster ($N_{U}>1$) to produce discriminative features for different unknown classes. Therefore, our approach first initializes a list of potential unknown clusters at each learning step via DB-SCAN [21] on the features of unknown classes extracted by the current CSS model. For the new known class $\mathcal{C}^{t}$, we initialize these clusters based on the mean of their feature representations. Meanwhile, the clusters of known classes learned in previous steps are maintained.
3.4 Continual Learning Procedure
Fig. 3 illustrates the training procedure of our continual learning approach. At each learning step $t$, the CSS model $F$ with $\theta_{t}$ is trained with the Fairness Contrastive Clustering loss defined in Eqn. (4) and the previous visual grammar model $\phi$ with $\Theta_{t-1}$. In addition, we introduce a cluster regularizer $\mathcal{R}_{C}$ to avoid the clusters of different classes collapsing into a single cluster. Therefore, the entire CSS learning objective in our approach can be formed as:
$\scriptsize\begin{split}\arg\min_{\theta_{t}}\mathbb{E}_{\mathbf{x}^{t},\mathbf{\hat{y}}^{t}}\Big{[}\mathcal{L}_{CE}\left(\mathbf{y}^{t},\mathbf{\hat{y}}^{t}\right)+\lambda_{CL}\sum_{\mathbf{c}_{i}}\mathcal{L}^{\alpha}_{Cont}\left(\mathbf{F}^{t},\mathbf{c}_{i}\right)+\lambda_{C}\mathcal{R}_{C}(\mathbf{c})\Big{]}\end{split}$ (8)
where $\mathcal{R}_{C}(\mathbf{c})=\sum_{\mathbf{c}_{i},\mathbf{c}_{j}}\{\max(0,2\nabla-||\mathbf{c}_{i}-\mathbf{c}_{j}||)\}^{2}$ is the regularizer to avoid the cluster collapsing, $\lambda_{C}$ is the balanced weight, and $\nabla$ is the margin between clusters.
Training Procedure of Visual Grammar Model At CSS learning step $t$, we adopt the visual grammar model trained on the previous learning step, i.e., $\phi$ with $\Theta_{t-1}$, to perform the cluster assignment for the contrastive clustering loss defined in Eqn. (2). Then, the visual grammar model at learning step $t$, i.e., $\phi$ with $\Theta_{t}$, will be learned (initialized from $\Theta_{t-1}$) on the features extracted from the dataset and the set of known clusters $\mathbf{c}$ up to the current learning step. Following [35], we sample a center $\mathbf{c}$ from the known clusters and its $M$ closest features to train the visual grammar model.
Initial Visual Grammar Model At the first learning step $t=1$, since no clusters have been learned at initial, the visual grammar model $\phi$ with $\Theta_{0}$ is not available. However, as common practices in CSS [19, 4, 5], the segmentation model is typically trained from a pre-trained backbone on ImageNet [17]. As a result, the features extracted at the first learning step are characterized by the ImageNet features. Therefore, we adopt this philosophy to initialize our visual grammar model ($\phi$ with $\Theta_{0}$) by pre-training the visual grammar model on the ImageNet dataset. Then, during CSS training, we will progressively train our visual grammar model at each learning step as aforementioned.
4 Experiments
4.1 Implementations and Evaluation Protocols
Table 1: Effectiveness of Fairness Contrastive Learning Loss.
(a) ADE20K 100-50
$\mathcal{L}_{Cont}$ $\mathcal{L}^{\alpha}_{Cont}$ 0-101 100-150 all Major Minor
44.6 15.2 34.8 51.5 26.4
44.6 24.5 37.9 52.1 30.8
(b) ADE20K 100-10
$\mathcal{L}_{Cont}$ $\mathcal{L}^{\alpha}_{Cont}$ 0-101 100-150 all Major Minor
41.9 16.0 33.2 49.9 24.9
44.4 20.4 36.4 51.8 28.7
Implementation Following common practices [19, 52, 3], we adopt DeepLab-V3 [6] with ResNet-101 [24] and SegFormer [56] with MiT-B3 [56] in our experiments. For the Visual Grammar model, we adopt the design of [35] with $L_{\phi}=12$ blocks of multi-head self-attention layers. The feature vectors from the last layer of the decoder are used for our $\mathcal{L}^{\alpha}_{Cont}$ loss. The value $\alpha$ is set individually for each dataset, i.e., $\alpha=5\times 10^{-2}$ for ADE20K, $\alpha=10^{-2}$ for VOC for Cityscapes. The details of our hyper-parameters are provided in the supplementary.
Table 2: Effectiveness of Scaling Factor $\alpha$ in $\mathcal{L}^{\alpha}_{Cont}$.
(a) ADE20K 100-50
$\alpha=0.1$ 43.1 19.8 35.3 50.6 27.7
$\alpha=0.05$ 44.6 24.5 37.9 52.1 30.8
$\alpha=0.01$ 43.6 21.3 36.2 51.0 28.7
$\alpha=0.005$ 42.4 18.6 34.5 50.1 26.6
(b) Pascal VOC 15-5
$\alpha$ 0-15 16-20 all Major Minor
$\alpha=0.1$ 74.8 51.6 69.3 76.9 63.5
$\alpha=0.05$ 76.2 51.3 70.3 79.0 63.8
$\alpha=0.01$ 79.4 54.8 73.5 81.3 67.7
$\alpha=0.005$ 74.6 48.9 68.5 77.6 61.6
Evaluation Protocols: We evaluate models on three standard datasets of CSS, i.e., ADE20K [61], Pascal VOC [22], and Cityscapes [15]. Following common practices [52, 4, 5], our experiments are conducted on the overlapped CSS settings. In particular, on ADE20K, we use three different settings, i.e., ADE20K 100-50 (2 steps), ADE20K 100-10 (6 steps), and ADE20K 100-5 (11 steps). On Pascal VOC, we evaluate FALCON in three benchmarks, i.e., VOC 15-5 (2 steps), VOC 15-1 (6 steps), and VOC 10-1 (11 steps). On Cityscapes, we conduct domain incremental experiments with three settings, i.e., Cityscapes 11-5 (3 steps), Cityscapes 11-1 (11 steps), and Cityscapes 1-1 (21 steps). Following [19, 4], the mean Intersection over Union (mIoU) metric is adopted in our comparison, including mIoU of the last learning step on initial classes, incremental classes, and all classes. In addition, to illustrate the fairness improvement, we report the mIoU of major and minor classes.
4.2 Ablation Study
Effectiveness of Fairness Contrastive Clustering Table 1 presents our results using DeepLab-V3 [7] with Resnet101 on ADE20K 100-50 and ADE20K 100-10 benchmarks. We evaluate the impact of the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}$ by comparing it with the vanilla contrastive clustering loss $\mathcal{L}_{Cont}$. As shown in our results, the overall performance has been significantly improved to $37.9\%$ and $36.4\%$ on ADE20K 100-50 and ADE20K 100-10, respectively. In addition, the fairness of the model has been promoted since the mIoU performance of major and minor groups was enhanced. We also study the impact of network backbones and cluster margin $\nabla$ in our supplementary.
Table 3: Effectiveness of Contributions of Our Proposed Losses.
(a) ADE20K 100-50
$\mathcal{L}_{CE}$ $\mathcal{L}^{\alpha}_{Cont}$ $\phi$ $\mathcal{R}_{C}$ 0-100 101-150 all Major Minor
0.0 18.9 6.3 0.0 9.4
44.0 7.9 31.9 51.6 22.1
43.8 21.8 36.4 51.1 29.1
44.6 24.5 37.9 52.1 30.8
(b) ADE20K 100-10
$\mathcal{L}_{CE}$ $\mathcal{L}^{\alpha}_{Cont}$ $\phi$ $\mathcal{R}_{C}$ 0-100 101-150 all Major Minor
0.0 3.5 1.2 0.0 1.8
39.0 13.1 30.4 47.8 21.6
43.4 18.5 35.1 51.2 27.1
44.4 20.4 36.4 51.8 28.7
Effectiveness of Scaling Factor of Cluster Table 2 illustrates the experimental results of the impact of different scaling factor $\alpha$ on ADE20K 100-50 and Pascal VOC 15-5 benchmarks. As shown in Table 2, when the value of scaling factor $\alpha$ gradually decreases, the performance of our proposed approach is improved accordingly since the fairness contrastive loss in Eqn (4) tends to be more uniform across major and minor classes. However, when the scaling factor is too small ($\alpha=0.005$), the impact of the loss enforcement becomes weaker leading to the weaker enforcement of the fairness contrastive clustering, resulting in lower overall performance. In addition, we have observed that the higher the number of classes demands the higher the value of $\alpha$ since it will increase the compactness of more clusters.
Figure 6: Cluster Distribution at Learning Step $t=1$ of ADE20K 100-50 (classes 109-144 are future classes) without (left) and with (right) Fairness Learning via Contrastive Attention.
Effectiveness of Loss Contributions Table 3 illustrates the contributions of proposed learning objectives. For the model without using visual grammar, we only use a single unknown cluster ($N_{U}=1$) and adopt the nearest cluster strategies to assign clusters of unknown pixels. By using only cross-entropy loss, the mIoU performance remains low due to catastrophic forgetting and background shift problems. Meanwhile, with our fairness clustering loss $\mathcal{L}^{\alpha}_{Cont}$, visual grammar model $\phi$, and the cluster regularizer $\mathcal{R}$, the mIoU performance has been significantly improved to $37.9\%$ and $36.4\%$ on ADE20K 100-50 and ADE20K 100-10, respectively. Moreover, our FALCON has significantly promoted the fairness of segmentation models illustrated by the mIoU improvement of major and minor groups.
Table 4: Effectiveness of Visual Grammar on ADE20K.
(a) ADE20K 100-50
0-100 101-150 all Major Minor
Nearest Cluster 44.3 11.5 33.4 51.5 24.3
Fixed $\phi$ 44.6 17.6 35.6 52.0 27.4
Adaptive $\phi$ 44.6 24.5 37.9 52.1 30.8
(b) ADE20K 100-10
0-100 101-150 all Major Minor
Nearest Cluster 40.1 14.3 31.5 48.7 22.9
Fixed $\phi$ 43.0 18.5 34.9 50.6 27.0
Adaptive $\phi$ 44.4 20.4 36.4 51.8 28.7
Table 5: Comparison with Prior Methods on ADE20K Benchmarks (Note: The results of MiB [3], PLOP [19], and FairCL [52] using Transformer on ADE20K 100-5 were not reported in prior studies. The upper bound results are not trained with fairness objective).
Network Method ADE20K 100-50 ADE20K 100-10 ADE20K 100-5
0-100 101-150 all avg 0-100 101-150 all avg 0-100 101-150 all avg
DeepLab-V3 PLOP [19] 41.9 14.9 32.9 37.4 40.5 14.1 31.6 36.6 39.1 7.8 28.8 35.3
RCIL [60] 42.3 18.8 34.5 -- 39.3 17.6 32.1 -- 38.5 11.5 29.6 --
REMINDER [39] 41.6 19.2 34.1 -- 39.0 21.3 33.1 -- 36.1 16.4 29.5 --
RCIL+LGKD [59] 43.3 25.1 37.2 -- 42.2 20.4 34.9 -- -- -- -- --
FairCL [52] 43.4 24.0 37.0 40.5 41.7 20.4 34.7 39.0 -- -- -- --
FALCON 44.6 24.5 37.9 41.3 44.4 20.4 36.4 40.1 38.0 16.1 30.7 37.6
[…]
MiB [3] 37.0 24.1 32.6 38.3 23.5 10.6 26.6 29.6 21.0 6.1 16.1 27.7
Mask2Former PLOP [19] 44.2 26.2 38.2 41.1 34.8 15.9 28.5 35.2 33.6 14.1 27.1 33.6
CoMFormer [4] 44.7 26.2 38.4 41.2 40.6 15.6 32.3 37.4 39.5 13.6 30.9 36.5
MiB [3] 43.4 30.6 39.2 38.7 39.1 20.4 34.2 39.5 -- -- -- --
PLOP [19] 43.8 26.2 38.0 38.1 43.3 24.1 36.2 40.3 -- -- -- --
Transformer FairCL [52] 43.6 25.5 37.6 40.7 42.2 21.9 35.5 39.4 -- -- -- --
FALCON 47.5 30.6 41.9 43.5 47.3 26.2 40.3 42.8 40.8 18.9 33.5 38.1
[…]
Table 6: Comparisons with Prior Methods on Pascal VOC.
Table 7: Comparisons on Cityscapes.
Method Pascal VOC 15-5 Pascal VOC 15-1 Pascal VOC 10-1
0-15 16-20 all 0-15 16-20 all 0-10 11-20 all
DeepLab-V3 MiB [3] 76.37 49.97 70.08 38.00 13.50 32.20 20.00 20.10 20.10
PLOP [19] 75.73 51.71 70.09 65.10 21.10 54.60 44.00 15.50 30.50
RCIL [60] -- -- -- 70.60 23.70 59.40 55.40 15.10 34.30
FairCL [52] -- -- -- 72.00 22.70 60.30 42.30 25.60 34.40
SSUL [5] 77.82 50.10 71.22 77.31 36.59 67.61 71.31 45.98 59.25
FALCON 79.35 54.77 73.50 78.34 42.57 69.83 73.94 49.73 62.41
[…]
Transformer PLOP [19] 72.51 48.37 66.76 64.59 37.23 58.08 48.53 33.71 41.47
SSUL [5] 79.91 56.83 74.41 79.91 40.56 70.54 74.06 51.85 63.48
FairCL [52] -- -- -- 73.50 22.80 61.50 57.10 14.20 36.60
FALCON 81.20 58.04 75.69 78.71 47.54 71.28 74.92 52.54 64.26
[…]
Method 11-5 11-1 1-1
DeepLab-V3 LWF-MC [42] 58.90 56.92 31.24
ILT [33] 59.14 57.75 30.11
MİB [3] 61.51 60.02 42.15
PLOP [19] 63.51 62.05 45.24
RCIL [60] 64.30 63.00 48.90
FairCL [52] 66.96 66.61 49.22
FALCON 70.74 69.75 55.24
[…]
Trans. FairCL [52] 67.85 67.09 55.68
FALCON 71.33 70.14 58.79
[…]
Table 7: Comparisons on Cityscapes.
Effectiveness of Visual Grammar We evaluate FALCON under three settings, i.e., Nearest Cluster, Fixed $\phi$ pretrained ImageNet (without updating on each learning step), and Adaptive $\phi$ (with updating on each learning step). As in Table 4, the mIoU result using only the nearest cluster remains ineffective. Meanwhile, the adaptive visual grammar model updated at each learning step further boosts the mIoU performance and promotes fairness, i.e., increased by $4.5\%$ and $4.9\%$ on ADE20K 100-50 and ADE20K 100-10 compared to the nearest cluster approach. In addition, we study the impact of choosing the number of features $M$ in the visual grammar model in our supplementary. Fig. 6 illustrates the feature distributions of unknown classes (future class). As a result, our FALCON approach is able to model features of unknown classes into different clusters and produce better and more compact clusters compared to the one without Fairness Learning via Contrastive Attention.
Figure 7: Visualization of Our Results on ADE20K 100-50.
4.3 Comparison with Prior SOTA Methods
ADE20K Table 5 presents our experimental results using DeepLab-V3 and Transformer networks compared to prior CSS methods. Overall, our proposed approach has achieved the SOTA performance compared to prior methods. In particular, by using DeepLab-V3, our approach has achieved SOTA performance, i.e., the mIoU results of $37.9\%$ and $+36.4\%$ on ADE20K 100-50 and ADE20K 100-10 benchmarks, higher than prior FairCL [52]. Meanwhile, our approach using Transformer has outperformed the prior SOTA CoMFormer [4] model by $+3.5\%$, $+8.0\%$, and $+2.6\%$ on ADE20K 100-50, ADE20K 100-10, and ADE20K 100-5 respectively. In addition, our mIoU results on the initial classes remain competitive with the upper-bounded results because our method is able to well handle the fairness problem compared to the fully supervised learning approach. We also report our results on the ADE20K 50-50 benchmark in the supplementary. As in Fig. 7, FALCON produces better segmentation maps compared to prior methods.
Pascal VOC Table 7 presents the results of our FALCON on Pascal VOC benchmarks. Our proposed approach has consistently achieved the SOTA performance on three benchmarks. In particular, compared to the prior FairCL [52] approach, our methods using DeepLab-V3 have improved the mIoU performance up to $73.50\%$, $69.83\%$, and $62.41\%$ on Pascal VOC 15-5, Pascal VOC 15-1, and Pascal VOC 10-1, respectively. Additionally, by using the better network backbone, i.e., Transformer, the performance of the segmentation model is also further improved. Our results have reduced the gap with the upper bound performance.
Cityscapes Table 7 reports the performance of our approach using DeepLab-V3 compared to prior methods on three different settings of Cityscapes benchmarks, i.e., Cityscapes 11-5, Cityscapes 11-1, and Cityscapes 1-1. As shown in the experimental results, the performance of our methods has consistently outperformed prior FairCL [52] approach by $+3.78\%$, $+3.14\%$, and $+6.02\%$ on three benchmarks. Similar to our experiments on ADE20K and VOC, the better network brings higher results. These results have shown the effectiveness of FALCON on various benchmarks.
5 Conclusions
This paper has presented a novel Fairness Learning via Contrastive Attention approach to CSS. In particular, the fairness contrastive clustering loss has been introduced to model both catastrophic forgetting and fairness problems. Then, the visual grammar model was presented to model the unknown classes. The experimental results on different benchmarks have shown the SOTA performance and fairness promotion of our proposed FALCON approach.
Limitations Our study chose a set of learning hyper-parameters to support our theoretical analysis. However, it potentially consists of several limitations related to choosing learning parameters and cluster initialization. The details of our limitations are discussed in the supplementary. These limitations will motivate future research to further improve our Fairness Learning via Contrastive Attention.
Supplementary
1 Proof of Propositions 1 and 2
1.1 Proof of Proposition 1
Proposition 1: If the contrastive clustering loss $\mathcal{L}_{Cont}(;,\mathbf{c})$ achieve the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converges to $\ell_{i}=L^{-1}$.
Proof: Let us consider the optimization of the Eqn. (4) in the paper as follows:
$\begin{split}&\min-\sum_{i=1}^{L}\log\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}=-\sum_{i=1}^{L}\log\ell_{i}\\ &\text{subject to}\quad\sum_{i=1}^{L}\ell_{i}=\ell\end{split}$ (9)
where $\ell$ is the total enforcement between features $\mathbf{f}^{t}_{i}$ and cluster $\mathbf{c}$. Then, the optimization of Eqn. (4) in the paper can be rewritten by using Lagrange multiplier as follows:
($\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)=-\sum_{i=1}^{L}\log\ell_{i}+\lambda(\sum_{i=1}^{L}\ell_{i}-\ell)$ (10)
where $\lambda$ is the Lagrange multiplier. Then, the contrastive clustering loss in Eqn. (4) in the paper achieves minimum if and only if:
($\begin{split}\frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\ell_{i}}&=-\ell_{i}^{-1}+\lambda=0\\ \frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\lambda}&=\sum_{i=1}^{L}\ell_{i}-\ell=0\\ \Rightarrow\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)&=-L\log\frac{\ell}{L}\end{split}$ (11)
As the total enforcement between features and the cluster is normalized, i.e., $\ell\in[0..1]$, the contrastive clustering loss ($\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)$ achieves minimum when $\log\ell=0\Rightarrow\ell=1$. Then, the enforcement between a single feature and the cluster will be equal to $\ell_{i}=\frac{\ell}{L}=L^{-1}$.
1.2 Proof of Proposition 2
Proposition 2: If the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}(;,\mathbf{c})$ achieve the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converges to $\ell_{i}=(\alpha^{-1}+L)^{-1}$.
Proof: We first define the the enforcement between transitive vector $\mathbf{v}$ and the cluster $\mathbf{c}$ as $\ell_{\mathbf{v}}=\frac{\exp(\mathbf{v}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}$. Then, let us consider the optimization of Eqn. (5) in the paper as follows:
$\begin{split}&\min-\sum_{i=1}^{L}\alpha\log\ell_{i}-\log\ell_{\mathbf{v}}\\ &\text{subject to}\quad\sum_{i=1}^{L}\ell_{i}+\ell_{\mathbf{v}}=\ell\end{split}$ (12)
Similar to Eqn. (9), Eqn. (12) can be reformulated via Lagrange multiplier as follows:
($\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)=-\sum_{i=1}^{L}\alpha\log\ell_{i}-\log\ell_{\mathbf{v}}+\lambda(\sum_{i=1}^{L}\ell_{i}+\ell_{\mathbf{v}}-\ell)$ (13)
Then, the fairness contrastive loss $\mathcal{L}^{\alpha}_{Cont}$ achieves minimum if and only if:
($\begin{split}\frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\ell_{i}}&=-\alpha\ell_{i}^{-1}+\lambda=0\\ \frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\ell_{\mathbf{v}}}&=-\ell_{\mathbf{v}}^{-1}+\lambda=0\\ \frac{\partial\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)}{\partial\lambda}&=\sum_{i=1}^{L}\ell_{i}+\ell_{\mathbf{v}}-\ell=0\\ \Rightarrow\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)&=-\alpha L\log\frac{\alpha\ell}{1+\alpha L}-\log\frac{\ell}{1+\alpha L}\end{split}$ (14)
As in Eqn. (14), the fairness contrastive learning loss ($\mathcal{L}\left(\{\ell_{i}\}_{i=1}^{L},\lambda\right)$ archives minimum when $\log\ell=0\rightarrow\ell=1$. Thus, the enforcement between the single feature the cluster will be re-balanced as $\ell_{i}=\frac{\alpha}{1+\alpha L}=(\alpha^{-1}+L)^{-1}$.
2 Implementation
Implementation Our framework is implemented in PyTorch and trained on four 40GB-VRAM NVIDIA A100 GPUs. The contrastive loss in our implementation is normalized with respect to the number of samples. These models are optimized by the SGD optimizer [2] with momentum 0.9, weight decay $10^{-4}$, and a batch size of $16$. The learning rate of the first learning step and the continual steps is set to $10^{-4}$ and $5\times 10^{-5}$ respectively. To update the cluster vectors $\mathbf{c}$, following prior work [52, 27, 25], we maintain a set of $500$ features for each cluster and update the clusters after $100$ steps with a momentum $\eta=0.99$. In our domain incremental experiments, all clusters are updated at each learning step by momentum update. The number of features selected for each cluster in the visual grammar model is set to $M=128$. The balanced weight of CSS objective $\lambda_{CL}$ and the cluster regularizer $\lambda_{C}$ is set to $1$. Following the common practices [52, 27], the margin between clusters $\nabla$ is set to 10.
Unknown Cluster Initialization As mentioned in the main paper, we adopt the DB-SCAN algorithm to initialize the clusters for unknown samples. In addition, to reduce the noise clusters and isolated clusters, we also merge several close clusters, i.e., if the distance between two clusters is less than the margin $2\nabla$, these will be merged into a single cluster where the new cluster center will be the means of these two merging cluster centers. By empirical observation, we have noticed that the number of unknown clusters initialized at each learning step, i.e., $N_{U}$ at the current learning step $t$, is not greater than 1.5$\times$ times of the remaining classes (i.e., $|\mathcal{C}^{t+1..T}|$) in the dataset, e.g., in our ADE20K 100-50 experiments, at the first learning step of $100$ classes, there are $68$ unknown clusters that have been initialized while there are $50$ remaining unknown classes in the dataset.
Cluster Assignment In our approach, we use our visual grammar model to assign the cluster for each feature representation. Theoretically, although there is a possibility that a feature could not be assigned to a cluster via the visual grammar model, we have empirically observed that this issue rarely happens in our approach. Indeed, since we initialize the known clusters via the DB-SCAN, it guarantees that for each feature, there is at least one cluster nearby that the feature representation should belong to. However, to preserve the integrity of our approach, for the outlier features in cases that cannot be assigned clusters via the visual grammar model, these outliers will be heuristically assigned to their closest clusters as similar to [27, 52].
Continual Learning Procedure Algorithm 1 illustrates the training procedure of our CSS approach.
Algorithm 1 CSS Procedure At Learning Step $t$
1: Learning Step $t$, Dataset $\mathcal{D}^{t}$, Visual Grammar $\phi(;\Theta_{t-1})$, and Segmentation Model $F(;\Theta_{t-1})$
2:Step 0: Extract features on $\mathcal{D}^{t}$ by $F(;,\theta_{t-1})$
3:Step 1: Initialize new known clusters for $\mathcal{C}^{t}$ of features extracted in Step 0
4:Step 2: Initialize potential unknown clusters of features extracted in Step 0
5:Step 3: Train CSS Model $F(;\theta_{t})$ on $\mathcal{D}^{t}$
6:Step 4: Extract features on $\mathcal{D}^{t}$ by $F(;,\theta_{t})$
7:Step 5: Train Visual Grammar Model $\phi(;,\Theta_{t})$ on current known clusters $\mathbf{c}$ and features extracted in Step 4
8:return $F(;\theta_{t})$ and $\phi(;\Theta_{t})$
3 Additional Experimental Results
3.1 Experiment Results of ADE20K 50-50 Benchmark
Table 8 presents the results of our method on the ADE20K 50-50 benchmark compared to prior methods. For fair comparisons, we use the DeepLab-V3 and Transformer in this experiment. As shown in the results, our proposed FALCON approach significantly outperforms prior methods. The results of our approach have reduced the gap with the upper bound result.
Table 8: Experimental results on ADE20K 50-50 Benchmark
ADE20K 50-50 (3 steps)
Network Method 0-50 50-150 all
MiB [3] 45.6 21.0 29.3
PLOP [19] 48.8 21.0 30.4
LGKD+PLOP [59] 49.4 29.4 36.0
DeepLab-V3 RCIL [60] 47.8 23.0 31.2
RCIL+LGKD [59] 49.1 27.2 34.4
FairCL [52] 49.7 26.8 34.6
FALCON 50.6 31.2 37.6
[…]
FairCL [52] 49.6 27.8 35.6
Transformer FALCON 53.0 36.8 42.2
[…]
3.2 Ablation Study
Effectiveness of Choosing Margin normal-$\nabla$ Table 9 studies the effectiveness of the value of margin $\nabla$ to the performance of our approach on ADE20K 100-50 and ADE20K 100-10 benchmarks. As shown in the results, the change of $\nabla$ also slightly influences the performance of the model. Since the margin defines the distance between two clusters, while the smaller value of the margin $\nabla$ could cause the incorrect cluster assignment of the features, the larger value of the margin $\nabla$ could produce the less compact clusters.
Table 9: Effectiveness of Choosing Margin $\nabla$
(a) ADE20K 100-50
0-100 101-150 all Major Minor
$\nabla=5$ 44.4 21.8 36.9 51.9 29.4
$\nabla=10$ 44.6 24.5 37.9 52.1 30.8
$\nabla=20$ 44.7 22.2 37.2 51.7 29.9
(b) ADE20K 100-10
0-100 101-150 all Major Minor
$\nabla=5$ 43.2 18.7 35.0 50.5 27.3
$\nabla=10$ 44.4 20.4 36.4 51.8 28.7
$\nabla=20$ 43.5 19.9 35.7 51.2 27.9
Effectiveness of Choosing Number of Features $M$ We study the impact of choosing the number of features $M$ in the visual grammar model. As in shown Table 10, the optimal performance of our approach is $M=128$. When the number of features selected is small ($M=96$), it does not have enough number of features to form the visual grammar so the model is hard to exploit the correlation among features and the cluster. Meanwhile, when we increase the number of selected features ($M=256$), the clusters will consist of many outlier features (the ones that do not belong to the cluster), thus being challenging for the visual grammar model to exploit the topological structures of the feature distribution.
Table 10: Effectiveness of Number of Features $M$ in a Cluster of Visual Grammar Model.
(a) ADE20K 100-50
0-100 101-150 all Major Minor
$M=96$ 43.0 19.6 35.2 50.5 27.5
$M=128$ 44.6 24.5 37.9 52.1 30.8
$M=256$ 43.6 21.6 36.3 51.0 28.9
(b) ADE20K 100-10
0-100 101-150 all Major Minor
$M=96$ 42.2 16.4 33.6 50.2 25.3
$M=128$ 44.4 20.4 36.4 51.8 28.7
$M=256$ 42.7 17.1 34.2 50.6 26.0
Effectiveness of Different Segmentation Networks To illustrate the flexibility of our proposed approach, we evaluate our proposed approach with different network backbones. Table 11 illustrates the results of our approach using DeepLab-V3 [7], SegFormer [56] with different backbones, i.e., ResNet-50, ResNet-101, MiT-B2, and MiT-B3. As shown in the performance, the more powerful the segmentation model is, the better performance of the model is. In particular, our approach has shown its flexibility since it consistently improves the performance of the segmentation model and achieves the SOTA performance on two different benchmarks, i.e., the performance of Transformer models achieves $41.9\%$, and $40.3\%$ on ADE20K 100-50, ADE20K 100-10, respectively.
Table 11: Effectiveness of Different Backbones on ADE20K.
(a) ADE20K 100-50
Backbone 0-100 101-150 all Major Minor
DeepLab-V3 R-50 44.3 15.2 34.7 51.5 26.4
R-101 44.6 24.5 37.9 52.1 30.8
[…]
MiT-B3 47.5 30.6 41.9 53.8 35.8
(b) ADE20K 100-10
Backbone 0-100 101-150 all Major Minor
DeepLab-V3 R-50 43.5 16.5 34.5 51.1 26.2
R-101 44.4 20.4 36.4 51.8 28.7
[…]
MiT-B3 47.3 26.2 40.3 54.0 33.4
4 Relation to Knowledge Distillation
Knowledge Distillation is a common approach to continual semantic segmentation [19, 4, 60, 5]. Prior work in clustering [52] has shown that the clustering loss is an upper bound of the knowledge distillation loss. Formally, the knowledge distillation loss can be formed as follows:
$\small\mathcal{L}_{distill}(\mathbf{x}^{t},F,\theta_{t},\theta_{t-1})=\mathcal{L}(\mathbf{F}^{t-1},\mathbf{F}^{t})$ (15)
where $\mathbf{F}^{t}$ and $\mathbf{F}^{t-1}$ are the feature representations extracted from the model at learning step $t$ and step $t-1$, respectively, and the metric $\mathcal{L}$ measure the knowledge gap between $\mathbf{F}^{t}$ and $\mathbf{F}^{t-1}$. Then, given a set of cluster $\mathbf{c}$, we consider the following triangle inequality of the metric $\mathcal{L}$ as follows:
$\small\begin{split}\forall\mathbf{c}:\quad\mathcal{L}(\mathbf{F}^{t},\mathbf{F}^{t-1})&\leq\mathcal{L}(\mathbf{F}^{t},\mathbf{c})+\mathcal{L}(\mathbf{c},\mathbf{F}^{t-1})\\ \Leftrightarrow\underbrace{\mathcal{L}(\mathbf{F}^{t},\mathbf{F}^{t-1})}_{\mathcal{L}_{distill}}&\leq\frac{1}{|\mathcal{C}^{1..T}|}\sum_{\mathbf{c}}\left[\underbrace{\mathcal{L}(\mathbf{F}^{t},\mathbf{c})}_{\mathcal{L}_{Cont}}+\mathcal{L}(\mathbf{c},\mathbf{F}^{t-1})\right]\end{split}$ (16)
At the computational time of Contrastive Clustering loss, the set of cluster vectors $\mathbf{c}$ is fixed (could be considered as constants). In addition, the features extracted at learning step $t-1$, i.e., $\mathbf{F}^{t-1}$, are constant due to the fix pre-trained model $\theta_{t-1}$. Therefore, without a strict argument, the distance $\mathcal{L}(\mathbf{c},\mathbf{F}^{t-1})$ could be considered as constant. Therefore, Eqn. (16) can be further derived as follows:
$\footnotesize\begin{split}\underbrace{\mathcal{L}(\mathbf{F}^{t},\mathbf{F}^{t-1})}_{\mathcal{L}_{distill}}&=\mathcal{O}\Bigg{(}\underbrace{\mathcal{L}\frac{1}{|\mathcal{C}^{1..T}|}}_{Constant}\sum_{\mathbf{c}}\Big{[}\underbrace{\mathcal{L}(\mathbf{F}^{t},\mathbf{c})}_{\mathcal{L}_{Cont}}+\underbrace{\mathcal{L}(\mathbf{c},\mathbf{F}^{t-1})}_{\text{Constant}}\Big{]}\Bigg{)}\\ &=\mathcal{O}\Bigg{(}\underbrace{\sum_{\mathbf{c}}\mathcal{L}(\mathbf{F}^{t},\mathbf{c})}_{\mathcal{L}_{Cont}}\Bigg{)}\\ \Rightarrow\mathcal{L}_{distill}(\mathbf{F}^{t-1},\mathbf{F}^{t})&=\mathcal{O}\left(\mathcal{L}_{Cont}(\mathbf{F}^{t},\mathbf{c})\right)\end{split}$ (17)
where $\mathcal{O}$ is the Big-O notation. Hence, from Eqn. (17), without lack of generality, we can observe that the Contrastive Clustering Loss is the upper bound of the Knowledge Distillation loss. Therefore, by minimizing the Contrastive Clustering Loss, the constraint of Knowledge Distillation is also maintained due to the property of the upper bound.
5 Discussion of Limitations
In our paper, we choose a specific set of hyper-parameters and learning approaches to support our hypothesis. However, our work could contain several limitations. First, choosing the scaling factor $\alpha$ could be considered as one of the potential limitations of our approach. In practice, when data keeps continuously growing, the pre-defined scaling factor $\alpha$ could not be good enough to control the fairness among classes. Our work focuses on investigating the effectiveness of our proposed losses to fairness, catastrophic forgetting, and background shift problems. Thus, the investigation of balance weights among losses has not been fully exploited, and we leave this experiment as our future work. Third, initializing the unknown clusters at each training step could potentially be room for improvement since the bad initial clusters could result in difficulty during training and updating these clusters and linking the unknown clusters learned in previous steps and new initial unknown clusters at the current learning steps have been yet fully exploited in our method. In addition, while our approach is designed for the DeepLab-V3 and Transformer segmentation networks [7, 56], the extensions of FALCON to mask-based segmentation networks [13, 12, 4] could be a potential next research for further performance improvement. These limitations could motivate new studies to further improve Fairness Learning via the Contrastive Attention Approach to continual learning in the future.