### v3

# FALCON  [Uncaptioned image]: Fairness Learning via Contrastive Attention Approach to Continual Semantic Scene Understanding

Abstract
Continual Learning in semantic scene segmentation aims to continually learn new unseen classes in dynamic environments while maintaining previously learned knowledge. Prior studies focused on modeling the catastrophic forgetting and background shift challenges in continual learning. However, fairness, another major challenge that causes unfair predictions leading to low performance among major and minor classes, still needs to be well addressed. In addition, prior methods have yet to model the unknown classes well, thus resulting in producing non-discriminative features among unknown classes. This work presents a novel Fairness Learning via Contrastive Attention Approach to continual learning in semantic scene understanding. In particular, we first introduce a new Fairness Contrastive Clustering loss to address the problems of catastrophic forgetting and fairness. Then, we propose an attention-based visual grammar approach to effectively model the background shift problem and unknown classes, producing better feature representations for different unknown classes. Through our experiments, our proposed approach achieves State-of-the-Art (SoTA) performance on different continual learning benchmarks, i.e., ADE20K, Cityscapes, and Pascal VOC. It promotes the fairness of the continual semantic segmentation model.
1 Introduction
Figure 1: Our Fairness Learning via Contrastive Attention to Continual Semantic Segmentation. The Fairness Contrastive Clustering Loss promotes the fairness of the model while the Attention-based Visual Grammar models the unknown classes.
The semantic segmentation networks, e.g., Transformers [70] and Convolutional Neural Networks [8], learned from data with a closed-set of known classes, have shown outstanding performance. However, they often suffer performance degradation when encountering novel objects or classes in new dynamic environments [21, 3, 63, 64]. To improve their performance, several transfer learning and domain adaptation methods [29, 1, 63, 60, 65, 58, 41, 61, 30, 59] were introduced to adapt trained models into deployed environments. While the former often aims to fine-tune the model on labeled data collected in the new environments, the latter adapts the model to the new domains in an unsupervised manner [1, 29]. However, these methods cannot handle novel objects well due to their close-set learning. In practice, the semantic segmentation models should be able to adaptively and continually learn the new knowledge of novel classes. It motivates the development of Continual Learning paradigm [3, 21, 47], a.k.a, Continual Semantic Segmentation (CSS), where the segmentation models are learned sequentially to new contents of data.
Far apart from prior segmentation methods [8, 70] that learn one time on static, closed-set data, Continual Learning requires the segmentation models to learn from dynamic, open-set data [21, 63]. In a particular scenario, accessing previous learning data is restricted due to privacy concerns. In CSS, three challenges have been identified, including (1) Catastrophic Forgetting, (2) Background Shift, and (3) Fairness. While the catastrophic forgetting problem [51, 25, 57] depicts the segmentation model tends to forget its knowledge when learning new data, background shift indicates the problem of classes of previous or future data (unknown classes) have collapsed into a background class [3, 21, 76]. Prior methods [21, 47, 5, 63] addressed these two problems by introducing knowledge distillation and pseudo labels. However, these methods can not handle unknown classes since they either consider these unknown classes as a background class or assign unknown pixels by a pseudo label of prior known classes [21, 63]. More importantly, the last problem, fairness, is a significant challenge that limits the performance of CSS models (Fig. 1).
Figure 2: The Data Class Distribution of ADE20K. The major classes occupy more than 75% of the total pixels of the dataset.
As shown in Fig. 2, the number of pixels of each class in training data have been imbalanced among classes and significantly decreased after each task. Thus, this bias influences the learning procedure and model predictions that later cause unfair class predictions. However, limited studies are taking the fairness problem into account. [62] presented a similar problem in domain adaptation and extended it to continual learning [63]. These methods rely on the assumption of ideal fair or balanced data distributions. However, it is not applicable in practice since the size, i.e., the number of pixels, of several classes can never be more significant than others. For example, the bottle size should not be more significant than the size of a car. Meanwhile, the current knowledge distillation methods [21, 46, 4] in CSS are unable to handle the fairness problem since they focus on modeling catastrophic forgetting and background shift problems. Therefore, it is essential to develop a new CSS approach to address these limitations.
Contributions of This Work. This work presents a novel Fairness Learning via Contrastive Attention Approach (FALCON) to Continual Semantic Segmentation (as shown in Fig. 1). Our contributions can be summarized as follows. First, we introduce a novel Contrastive Clustering Paradigm approach to Continual Learning that models the catastrophic forgetting problem. Second, by analyzing the limitation of vanilla Contrastive Clustering in biased data, we introduce a novel Fairness Contrastive Clustering loss to model the fairness problem in continual learning efficiently. Third, to effectively model the background shift problem, we introduce a new Attention-based Visual Grammar that model the topological structures of feature distribution to handle the unknown classes effectively. Finally, the ablation studies illustrate the effectiveness of the proposed approach in different aspects of fairness promotion in CSS models. Compared with prior methods, our approach achieves SoTA performance on different settings of three standard benchmarks of CSS, including ADE20K, Pascal VOC, and Cityscapes.
2 Related Work
Continual Semantic Segmentation. Several studies were introduced to address catastrophic forgetting and background shift problems [22, 33, 38, 20, 21, 45, 44, 40]. The common CSS approach adopts knowledge distillation [21] and pseudo labels [3] to model catastrophic forgetting and background shift, respectively. Later, it was further improved by decoupling knowledge representations [76], modeling the inter- and intra-class knowledge [47, 16], distinguishing the feature representations of the future classes [5, 46, 68, 71], reducing background confusion [74, 75], using visual prompting [32, 36], or modeling distillation loss via the geodesic flow [55]. Another approach [4, 26, 6] adopted the mask-based segmentation networks [13, 14] to improve the performance. Recent studies have introduced CSS under the unsupervised domain adaptation settings [67, 52, 53, 64]. However, prior studies have yet to well model different unknown classes. Particularly, as previous methods [21, 4] use the pseudo labels to model the unknown classes, the future classes will be treated as a single background class. Then, Joseph et al. [31] improves the unknown class modeling using clustering but this method considers all different unknown classes as a single cluster, leading to the non-discriminative features among unknown classes.
Contrastive Learning (CL). CL is a common learning approach [43, 11, 10] to structure the deep feature representations in the deep latent space. Oorde et al. [43] first introduced the Noise-Contrastive Estimation (InfoNCE) learning framework. Then, Chen et al. [10] presented SimCLR, a self-supervised contrastive learning approach to improve the representation power of Residual Networks. He [28] proposed a Momentum Contrast framework for unsupervised representation learning. Later, it was further improved by using MLP projection head [11] and extended to improve the self-supervised training process of vision transformers [12]. Cui et al. [18] introduced a supervised parametric contrastive learning loss to address the long-tailed recognition. Li et al. [35] adopted contrastive learning to develop the one-stage online contrastive clustering method. Radford et al. [48] presents a contrastive framework to learn the vision-language model. Later, several methods also adopted this framework for vision-language pretraining [48, 34].
Imbalanced and Fairness Learning. The early methods utilized the balanced Softmax loss [50] to alleviate the impact of imbalanced data distribution. Later, Wang et al. [69] introduced a Seesaw loss to re-balance the contributions of positive and negative instances via the mitigation and compensation modules. Ziwei et al. [37] introduced a dynamic meta-embedding to model the imbalanced classification problem. Chu et al. [15] reduce the bias in the segmentation model by presenting a new stochastic training scheme. Szabo et al. [56] presented a tilted cross-entropy loss to promote class-relevant fairness. However, there are limited studies that address the fairness problem in CSS. Truong et al. [62] introduced a fairness domain adaptation approach to semantic segmentation and later extended it into continual learning setting [63]. However, these methods [62, 63] rely on the assumption of ideal balanced data, which could not be achieved by nature. To address the limitations in prior work, this paper will introduce a novel approach to effectively model the fairness problem and unknown classes in the continual learning setting.
Figure 3: The Proposed Fairness Learning via Contrastive Attention Approach to Continual Semantic Scene Understanding.
3 The Proposed FALCON Approach
CSS aims to learn a segmentation network $F$ on sequence data $\mathcal{D}=\{\mathcal{D}^{1},...,\mathcal{D}^{T}\}$ where $T$ is the number of learning steps. At learning step $t$, the model $F$ encounters a dataset $\mathcal{D}_{t}=\left\{(\mathbf{x}^{t},\mathbf{\hat{y}}^{t})\right\}$ where $\mathbf{x}^{t}\in\mathbb{R}^{H\times W\times 3}$ is the image and $\mathbf{y}\in\mathbb{R}^{H\times W}$ is a segmentation label of $\mathbf{x}^{t}$. The ground truths at learning step $t$ only consist of current classes $\mathcal{C}^{t}$, while the class labels of the previous $\mathcal{C}^{1...t-1}$ and future steps $\mathcal{C}^{t+1...T}$ are collapsed into a background class. Formally, learning the CSS model at step $t$ can be formed as Eqn. (1).
$\footnotesize\theta_{t}^{*}=\arg\min_{\theta_{t}}\mathbb{E}_{\mathbf{x}^{t},\mathbf{\hat{y}}^{t}\in\mathcal{D}^{t}}\left[\mathcal{L}_{CE}\left(\mathbf{y}^{t},\mathbf{\hat{y}}^{t}\right)+\lambda_{CL}\mathcal{L}_{CL}\left(F(\mathbf{x}^{t})\right)\right]$ (1)
where, $\mathbf{y}^{t}=F(\mathbf{x}^{t},\theta_{t})$, $\theta_{t}$ is the parameter of $F$ at current learning step $t$, $\mathcal{L}_{CE}$ is the cross-entropy loss, $\lambda_{CL}$ is the balanced weight. and $\mathcal{L}_{CL}$ is the CSS objective. At learning step $t$, the segmentation model $F$ is required to be able to predict both previously learned classes $\mathcal{C}^{1...t-1}$ and current new classes $\mathcal{C}^{t}$. Under this learning scenario, three challenges have been identified, i.e., Catastrophic Forgetting, Background Shift, and Fairness. Several prior methods were presented to model the two first issues in CSS using knowledge distillation [21, 4]. The last issue has not been well addressed yet due to its challenges [63]. Prior methods [54, 3, 21, 76, 4] adopt knowledge distillation to design $\mathcal{L}_{CL}$. However, this method prevents the CSS model from diverging knowledge learned previously, therefore resulting in limiting the ability to adopt new knowledge [63]. In addition, these methods have not addressed fairness and background shift problems due to their dedicated design for maintaining knowledge via distillation [21, 4, 46]. Therefore, to address these problems, we introduce a novel Fairness Learning via Contrastive Attention Approach to CSS.
3.1 Continual Learning via Contrastive Clustering
Apart from prior methods [21, 4, 46], our CSS (Fig. 3) is defined as Contrastive Clustering Learning. Given a set of centroid vectors $\{\mathbf{c}_{i}\}_{i=1}^{N_{K}+N_{U}}$ where $N_{K}=|\mathcal{C}^{1..t}|$ and $N_{U}$ is the number of known and unknown classes up to current learning tasks. Prior work [31, 63, 4] often defined the number of unknown classes as $1$ where background classes are considered as a single unknown class. Formally, our Contrastive Clustering Learning for CSS can be defined as in Eqn. (2).
$\footnotesize\begin{split}\mathcal{L}_{CL}\left(F(\mathbf{x}^{t})\right)&=\sum_{\mathbf{c}_{i}}\mathcal{L}_{Cont}(\mathbf{F}^{t},\mathbf{c}_{i})\\ &=\sum_{\mathbf{c}_{i}}\sum_{h,w}-\phi(\mathbf{f}^{t}_{h,w},\mathbf{c}_{i})\log\frac{\exp(\mathbf{f}^{t}_{h,w}\times\mathbf{c}_{i})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c}_{i})}\end{split}$ (2)
where $\mathbf{F}^{t}\in\mathbb{R}^{H\times W\times D}$ is the feature maps extracted from the input image $\mathbf{x}^{t}$ by the segmentation network $F$, $\mathbf{f}^{t}_{h,w}\in\mathbb{R}^{D}$ is the feature at the pixel location $(h,w)$ of features $\mathbf{F}^{t}$, $\sum_{\mathbf{f}^{\prime}}$ means the summation over all feature representations $\mathbf{f}^{\prime}\in\mathbb{R}^{D}$, and $\phi:\mathbb{R}^{D}\times\mathbb{R}^{D}\to[0,1]$ is the function that determines either $\mathbf{f}^{t}_{h,w}$ belongs to the cluster $\mathbf{c}_{i}$ or not.
By defining CSS as contrastive clustering learning, the knowledge of the segmentation model has been well maintained via the cluster vectors $\mathbf{c}$. Then, minimizing Eqn. (2) will separate the representations of different classes while gathering the features of the same class into the same cluster. As the cluster vectors $\mathbf{c}$ of the old classes $\mathcal{C}^{1..t-1}$ have been well learned to represent their knowledge, these vectors are frozen at learning step $t$ to maintain the knowledge representations of previous classes to address the catastrophic forgetting problem. To effectively learn cluster vectors $\mathbf{c}$, the cluster vector $\mathbf{c}$ will periodically updated after each $K$ steps by the momentum update [31, 63, 28] based on the features $\mathbf{f}^{t}_{h,w}$ assigned to cluster $\mathbf{c}$. However, there are two major problems in contrastive clustering learning. First, since the training data suffer the bias among classes as shown in Fig. 2, this bias will influence Eqn. (2) and cause the unfair predictions. Second, as the function $\phi$ requires the labels to determine the features belonging to clusters, it limits the ability to model the unknown classes where the labels are unavailable. Therefore, Secs. 3.2-3.3 will present a novel approach to tackle these problems.
3.2 Fairness Contrastive Clustering Learning
While contrastive clustering learning in Eqn. (2) promotes the compact representations of features around their clusters, inspired by [78, 18, 9], we observe that the imbalanced class distribution will influence unfair behaviors among classes. In particular, for simplicity, we consider $\{\mathbf{f}^{t}_{i}\}_{i=1}^{L}$ is the set of features that belong to the cluster $\mathbf{c}$ at learning step $t$ (i.e., $\phi(\mathbf{f}_{i}^{t},\mathbf{c})=1$) and $L$ is the number of features (in this case, $L$ is the total number of pixels belong to the class of cluster $\mathbf{c}$). Let us define the enforcement between the feature $\mathbf{f}^{t}_{t}$ and the cluster $\mathbf{c}$ as $\ell_{i}=\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}$. Hence, the lower the value of the enforcement $\ell_{i}$ is, the more compact the representation of visual features and clusters is. Then, the contrastive clustering learning loss in Eqn. (2) of entire cluster $\mathbf{c}$ can be defined as in Eqn. (3).
$\small\mathcal{L}_{Cont}(;,\mathbf{c})=-\sum_{i=1}^{L}\log\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}=-\sum_{i=1}^{L}\log\ell_{i}$ (3)
Proposition 1. If the contrastive clustering loss $\mathcal{L}_{Cont}(;,\mathbf{c})$ achieves the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converge to $\ell_{i}=L^{-1}$.
Proposition 1 has implied that the class with more samples will result in a lower value of the enforcement and produce a more compact representation while the class having fewer samples will be more scattered in the feature space due to the higher value of the enforcement. In particular, let $L_{major}$ and $L_{minor}$ be the number of samples of the major and minor class where $L_{major}>>L_{minor}$. Then, based on Proposition 1, the enforcement between features and the cluster of the major class will be significantly lower than the one of the minor class, i.e., $L_{major}^{-1}<<L_{minor}^{-1}$. Therefore, a direct adoption of the contrastive clustering loss in Eqn. (2) will result in an unfair CSS model. In addition, for classes in the minority group, the lower value of $L$ results in the feature presentations of classes being far away from their clusters. Thus, the model will produce non-discriminative features compared to the ones in the majority group. Moreover, if the loss is applied to the cases of unknown labels, these feature representations can be scattered in the latent space and pulled into the incorrect clusters due to weak enforcement between features and clusters (Fig. 4).
To address the unfair problem in contrastive clustering learning, inspired by [78, 18, 9], we introduce a scaling factor $\alpha$ and a learnable transition vector $\mathbf{v}$ for each cluster $\mathbf{c}$ (all clusters have the same value of $\alpha$ but different vector $\mathbf{v}$). Our Fairness Contrastive Clustering Learning Loss for the entire cluster in Eqn. (3) can be re-formed as follows,
$\footnotesize\mathcal{L}^{\alpha}_{Cont}(;,\mathbf{c})=-\alpha\sum_{i=1}^{L}\log\frac{\exp(\mathbf{f}^{t}_{i}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}-\log\frac{\exp(\mathbf{v}\times\mathbf{c})}{\sum_{\mathbf{f}^{\prime}}\exp(\mathbf{f}^{\prime}\times\mathbf{c})}$ (4)
Intuitively, the scaling factor $\alpha$ will help to re-scale the impact of the enforcement in learning, and the transitive vector $\mathbf{v}$ assists in translating the center cluster into the proper position of the latent space. This action promotes the compactness of clusters in the minority group.
Proposition 2. If the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}(;,\mathbf{c})$ achieves the optimal value, the enforcement $\ell_{i}$ between the feature and the cluster will converge to $\ell_{i}=(\alpha^{-1}+L)^{-1}$. Proofs of Propositions 1-2 are in the appendix.
Figure 4: The Enforcement Loss of Contrastive Clustering $\mathcal{L}_{Cont}$ and Fairness Contrastive Clustering $\mathcal{L}^{\alpha}_{Cont}$ on Pascal VOC. Since $\mathcal{L}_{Cont}$ suffers severe biased, its clusters of minor classes remain scattered. Our $\mathcal{L}^{\alpha}_{Cont}$ produces a more uniform loss among classes, promotes fairness and compactness of clusters.
Under the Proposition 2, when the value of $\alpha$ is small, the divergence of the enforcement between major and minor classes will be smaller, i.e., $||(\alpha^{-1}+L_{major})^{-1}-(\alpha^{-1}+L_{minor})^{-1}||<||L^{-1}_{major}-L^{-1}_{minor}||$. Fig. 4 has illustrated the impact of fairness contrastive clustering loss. Therefore, our designed proposed fairness contrastive loss has effectively addressed the fairness issue in Eqn. (2). It should be noted that although the smaller $\alpha$ results in the fairer enforcement varied from major to minor classes. However, if the value of scaling factor $\alpha$ is too small, the contrastive clustering loss will rely more on the enforcement of the transitive vector $\mathbf{v}$, and the distribution of features $\mathbf{f}^{t}_{i}$ around its cluster $\mathbf{c}$ will be scattered due the weak enforcement caused by small $\alpha$. Therefore, the value of scaling factor $\alpha$ in practice should be carefully selected.
3.3 An Efficient Unknown Class Modeling
An ideal CSS approach must be able to model the unknown classes without supervision, especially in open-set contexts [31, 63] where there could be multiple unknown classes or objects. Prior studies have adopted the pseudo-label strategies [21, 4] based on the model predictions to assign labels for seen classes while unseen classes have been ignored, thus resulting in non-discriminative features. [31, 63] improved the background modeling by using an additional prototypical representation for unknown classes. However, these approaches consider different unknown classes as one (i.e., $N_{U}=1$) resulting in non-distinguished representations of different unknown classes. Thus, modeling function $\phi$ in Eqn. (2) without supervision of different unknown classes (i.e., $N_{U}>1$) is challenging.
Although modeling $\phi$ to determine the single feature $\mathbf{f}$ belonging to the cluster $\mathbf{c}$ is challenging, prior studies in clustering [42, 73, 72] have suggested that determine a set of features $\{\mathbf{f}^{t}_{i}\}_{i=1}^{M}$ belonging to cluster $\mathbf{c}$ should be easier.
Figure 5: The Proposed Visual Grammar Model.
This derives from the fact that even though the feature representations of different classes are different, the distributions of features around its cluster (termed as Visual Grammar) in the feature space should be similar among classes or clusters. As a result, by learning the distribution of features and their clusters, the model $\phi$ can determine whether a feature belongs to a cluster. Then, by learning the model $\phi$ on prior known clusters and features, the knowledge of $\phi$ can be adaptively applied to unknown clusters. Fig. 5 illustrates our visual grammar model of the cluster distributions.
Limitations of Prior Clustering Methods. The traditional methods in clustering, e.g., KNN or density-based clustering [23], remain limited to noisy features leading to producing the incorrect cluster assignment. Meanwhile, the modern clustering methods, e.g., Graph Neural Networks (GNNs) [73, 72], require a large memory to build the affinity graph for clusters. In addition, GNNs often learn the local structures of graphs (or clusters) and accumulate them via the aggregation layers. Hence, the global structures of the clusters, i.e., visual grammar, are not well modeled by GNNs [42]. Therefore, to address these limitations, we introduced a new Attention-based Visual Grammar approach to efficiently model the distribution of features and their clusters via the self-attention mechanism [66].
Remark 1: Given a center $\mathbf{c}$ and a set of $M$ features $\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}$ where $\mathbf{f}^{\mathbf{c}}_{i}$ denotes the feature $\mathbf{f}_{i}$ belonging to the cluster $\mathbf{c}$, and $\forall i\in[1..M-1]:\cos(\mathbf{f}_{i}^{\mathbf{c}},\mathbf{c})\geq\cos(\mathbf{f}_{i+1}^{\mathbf{c}},\mathbf{c})$ the Visual Grammar of the cluster $\mathbf{c}$ parameterized by $\Theta$ can be defined as in Eqn. (5).
$\footnotesize\begin{split}\Theta^{*}&=\arg\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{f}^{\mathbf{c}}_{1},\mathbf{f}^{\mathbf{c}}_{2},...,\mathbf{f}^{\mathbf{c}}_{M},\mathbf{c},\Theta)\right]\\ \Leftrightarrow\Theta^{*}&=\arg\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{\Delta}^{\mathbf{c}}_{1},\mathbf{\Delta}^{\mathbf{c}}_{2},...,\mathbf{\Delta}^{\mathbf{c}}_{M},\mathbf{c},\Theta)\right]\end{split}$ (5)
where $\Delta^{\mathbf{c}}_{i}=\mathbf{f}^{\mathbf{c}}_{i}-\mathbf{c}$. Eqn. (5) defines the visual grammar of the cluster by modeling the feature distribution of $\mathbf{f}_{i}^{\mathbf{c}}$ and its cluster center $\mathbf{c}$. Let $\phi:\mathbb{R}^{(M+1)\times D}\to[0,1]^{M}$ be a function receiving a center $\mathbf{c}$ and a set of $M$ features $\{\mathbf{f}_{i}\}_{i=1}^{M}$ $\left(\cos(\mathbf{f}_{i},\mathbf{c})\geq\cos(\mathbf{f}_{i+1},\mathbf{c})\right)$ to determine whether $\mathbf{f}_{i}$ belonging to $\mathbf{c}$, i.e., $\mathbf{u}=\phi(\mathbf{\Delta}_{1},\mathbf{\Delta}_{2},...,\mathbf{\Delta}_{M},\mathbf{c})$ where $\mathbf{\Delta}_{i}=\mathbf{f}_{i}-\mathbf{c}$, $\mathbf{u}=[u_{1},u_{2},...,u_{M}]$ and $u_{i}=1$ denotes $\mathbf{f}_{i}$ belong to cluster $\mathbf{c}$ and vice versa. Hence, the visual grammar model in Eqn. (5) can be modeled by the network $\phi$ as follows with parameter $\Theta$ as in Eqn. (6).
$\footnotesize\begin{split}\Theta^{*}&=\arg\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{u}|\mathbf{\Delta}_{1},\mathbf{\Delta}_{2},...,\mathbf{\Delta}_{M},\mathbf{c},\Theta)\right]\end{split}$ (6)
Eqn. (6) aims to model the distribution of features around its cluster by learning the correlation of relatively topological structures $\mathbf{\Delta}_{i}$ of features $\mathbf{f}_{i}$ around cluster $\mathbf{c}$. Then, based on knowledge of the cluster distribution, the model $\phi$ is able to determine whether a feature $\mathbf{f}_{i}$ belongs to cluster $\mathbf{c}$. Hence, it is essential that the model $\phi$ has the ability to exploit the correlation between features $\mathbf{f}_{i}$ and cluster $\mathbf{c}$ to learn the topological structure of visual grammar. Therefore, we adopt the self-attention mechanism [66, 42] to efficiently model these feature correlations. Particularly, the model $\phi$ is formed by $L_{\phi}$ blocks of self-attention as follows,
$\footnotesize\begin{split}\mathbf{z}_{0}&=\operatorname{LN([\mathbf{\Delta}_{1},...,\mathbf{\Delta}_{M},\mathbf{c}])}+\boldsymbol{\beta},\quad\mathbf{a}_{l}=\mathbf{z}_{l}+\operatorname{MHSA}(\mathbf{z})\\ \mathbf{z}_{l+1}&=\mathbf{a}_{l}+\operatorname{MLP}(\operatorname{LN}(\mathbf{a}_{l})),\quad\quad\quad\quad\mathbf{u}=\operatorname{Proj}(\mathbf{z}_{L_{\phi}})\end{split}$ (7)
where $\boldsymbol{\beta}$ is the positional embedding, $\operatorname{LN}$ is Layer Normalization, $\operatorname{MHSA}$ is multi-head self-attention, $\operatorname{MLP}$ is the multi-layer perception, and $\operatorname{Proj}$ is the linear projection. By using Transformers, the correlation of cluster distributions can be well modeled by the self-attention mechanism.
Cluster Assignment via Visual Grammar. Instead of assigning the clusters based on the model prediction [21, 4, 46] or nearest cluster [63, 31] that are less effective, the cluster assignment in our approach will be performed by the visual grammar model, i.e., the visual grammar model will consider the $M$ closest features around cluster $\mathbf{c}$ to assign the cluster for these features. Then, the cluster assignments are used to compute our Fairness Contrastive Clustering loss. In addition, following common practices [21, 4, 46], we improve background shift modeling by using the cluster assignments of features as the pseudo labels of pixels.
Unknown Cluster Initialization. Prior work [63, 31] initialized a single unknown cluster ($N_{U}=1$), thus resulting in producing non-discriminative class-wise features. However, there should be more than a single unknown cluster ($N_{U}>1$) to produce discriminative features for different unknown classes. Therefore, our approach first initializes a list of potential unknown clusters at each learning step via DB-SCAN [23] on the features of unknown classes extracted by the current CSS model. For the new known class $\mathcal{C}^{t}$, we initialize these clusters based on the mean of their feature representations. Meanwhile, the clusters of known classes learned in previous steps are maintained. Further details can be found in the appendix.
3.4 Continual Learning Procedure
Fig. 3 illustrates the training procedure of our continual learning approach. At each learning step $t$, the CSS model $F$ with $\theta_{t}$ is trained with the Fairness Contrastive Clustering loss defined in Eqn. (4) and the previous visual grammar model $\phi$ with $\Theta_{t-1}$. In addition, we introduce a cluster regularizer $\mathcal{R}_{C}$ to avoid the clusters of different classes collapsing into a single cluster. Therefore, the entire CSS learning objective in our approach can be formed as in Eqn. (8).
$\scriptsize\vspace{-3mm}\begin{split}\arg\min_{\theta_{t}}\mathbb{E}_{\mathbf{x}^{t},\mathbf{\hat{y}}^{t}}\Big{[}\mathcal{L}_{CE}\left(\mathbf{y}^{t},\mathbf{\hat{y}}^{t}\right)+\lambda_{CL}\sum_{\mathbf{c}_{i}}\mathcal{L}^{\alpha}_{Cont}\left(\mathbf{F}^{t},\mathbf{c}_{i}\right)+\lambda_{C}\mathcal{R}_{C}(\mathbf{c})\Big{]}\end{split}$ (8)
where $\mathcal{R}_{C}(\mathbf{c})=\sum_{\mathbf{c}_{i},\mathbf{c}_{j}}\{\max(0,2\nabla-||\mathbf{c}_{i}-\mathbf{c}_{j}||)\}^{2}$ is the regularizer to avoid the cluster collapsing, $\lambda_{C}$ is the balanced weight, and $\nabla$ is the margin between clusters.
Training Procedure of Visual Grammar Model. At CSS learning step $t$, we adopt the visual grammar model trained on the previous learning step, i.e., $\phi$ with $\Theta_{t-1}$, to perform the cluster assignment for the contrastive clustering loss defined in Eqn. (2). Then, the visual grammar model $\phi$ with $\Theta_{t}$ at learning step $t$ will be learned (initialized from $\Theta_{t-1}$) on the features extracted from the dataset and the set of known clusters $\mathbf{c}$ up to the current learning step. Following [42], we sample a center $\mathbf{c}$ from the known clusters and its $M$ closest features to train the visual grammar model.
Initial Visual Grammar Model. At the first learning step $t=1$, since no clusters have been learned initially, the visual grammar model $\phi$ with $\Theta_{0}$ is unavailable. However, as common practices in CSS [21, 4, 5], the segmentation model is typically trained from a pre-trained backbone on ImageNet [19]. As a result, the features extracted at the first learning step are characterized by the ImageNet features. Therefore, we adopt this philosophy to initialize our visual grammar model ($\phi$ with $\Theta_{0}$) by pre-training the visual grammar model on the ImageNet dataset. Then, during CSS training, we will progressively train our visual grammar model at each learning step as aforementioned.
4 Experiments
4.1 Implementations and Evaluation Protocols
Table 1: Effectiveness of Fairness Contrastive Learning Loss.
(a) ADE20K 100-50
$\mathcal{L}_{Cont}$ $\mathcal{L}^{\alpha}_{Cont}$ 0-100 100-150 all Major Minor
44.6 15.2 34.8 51.5 26.4
44.6 24.5 37.9 52.1 30.8
(b) ADE20K 100-10
$\mathcal{L}_{Cont}$ $\mathcal{L}^{\alpha}_{Cont}$ 0-100 100-150 all Major Minor
41.9 16.0 33.2 49.9 24.9
44.4 20.4 36.4 51.8 28.7
Table 2: Effectiveness of Visual Grammar.
(a) ADE20K 100-50
0-100 101-150 all Major Minor
Nearest Cluster 44.3 11.5 33.4 51.5 24.3
Fixed $\phi$ 44.6 17.6 35.6 52.0 27.4
Adaptive $\phi$ 44.6 24.5 37.9 52.1 30.8
(b) ADE20K 100-10
0-100 101-150 all Major Minor
Nearest Cluster 40.1 14.3 31.5 48.7 22.9
Fixed $\phi$ 43.0 18.5 34.9 50.6 27.0
Adaptive $\phi$ 44.4 20.4 36.4 51.8 28.7
Implementation. Following common practices [21, 63, 3], we adopt DeepLab-V3 [7] with ResNet-101 [27] and SegFormer [70] with MiT-B3 [70] in our experiments. For the Visual Grammar model, we adopt the design of [42] with $L_{\phi}=12$ blocks of multi-head self-attention layers. The feature vectors from the last layer of the decoder are used for our $\mathcal{L}^{\alpha}_{Cont}$ loss. The value $\alpha$ is set individually for each dataset, i.e., $\alpha=5\times 10^{-2}$ for ADE20K, $\alpha=10^{-2}$ for VOC for Cityscapes. The details of our hyper-parameters are provided in the appendix.
Table 3: Effectiveness of Scaling Factor $\alpha$.
(a) ADE20K 100-50
$\alpha$ 0-15 16-20 all Major Minor
$\alpha=0.1$ 43.1 19.8 35.3 50.6 27.7
$\alpha=0.05$ 44.6 24.5 37.9 52.1 30.8
$\alpha=0.01$ 43.6 21.3 36.2 51.0 28.7
$\alpha=0.005$ 42.4 18.6 34.5 50.1 26.6
(b) Pascal VOC 15-5
$\alpha$ 0-15 16-20 all Major Minor
$\alpha=0.1$ 74.8 51.6 69.3 76.9 63.5
$\alpha=0.05$ 76.2 51.3 70.3 79.0 63.8
$\alpha=0.01$ 79.4 54.8 73.5 81.3 67.7
$\alpha=0.005$ 74.6 48.9 68.5 77.6 61.6
Table 4: Effectiveness of Our Proposed Losses.
(a) ADE20K 100-50
$\mathcal{L}_{CE}$ $\mathcal{L}^{\alpha}_{Cont}$ $\phi$ $\mathcal{R}_{C}$ 0-100 101-150 all Major Minor
0.0 18.9 6.3 0.0 9.4
44.0 7.9 31.9 51.6 22.1
43.8 21.8 36.4 51.1 29.1
44.6 24.5 37.9 52.1 30.8
(b) ADE20K 100-10
$\mathcal{L}_{CE}$ $\mathcal{L}^{\alpha}_{Cont}$ $\phi$ $\mathcal{R}_{C}$ 0-100 101-150 all Major Minor
0.0 3.5 1.2 0.0 1.8
39.0 13.1 30.4 47.8 21.6
43.4 18.5 35.1 51.2 27.1
44.4 20.4 36.4 51.8 28.7
Evaluation Protocols. We evaluate models on three standard datasets of CSS, i.e., ADE20K [77], Pascal VOC [24], and Cityscapes [17]. Following common practices [63, 4, 5], our experiments are conducted on the overlapped CSS settings. In particular, on ADE20K, we use three different settings, i.e., ADE20K 100-50 (2 steps), ADE20K 100-10 (6 steps), and ADE20K 100-5 (11 steps). On Pascal VOC, we evaluate FALCON in three benchmarks, i.e., VOC 15-5 (2 steps), VOC 15-1 (6 steps), and VOC 10-1 (11 steps). On Cityscapes, we conduct domain incremental experiments with three settings, i.e., Cityscapes 11-5 (3 steps), Cityscapes 11-1 (11 steps), and Cityscapes 1-1 (21 steps). Following [21, 4], the mean Intersection over Union (mIoU) metric is adopted in our comparison, including mIoU of the last learning step on initial classes, incremental classes, and all classes. In addition, to illustrate the fairness improvement, we report the mIoU of major and minor classes.
4.2 Ablation Study
Effectiveness of Fairness Contrastive Clustering. Table 1 presents our results using DeepLab-V3 [8] with Resnet101 on ADE20K 100-50 and ADE20K 100-10 benchmarks. We evaluate the impact of the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}$ by comparing it with the vanilla contrastive clustering loss $\mathcal{L}_{Cont}$. As shown in our results, the overall performance has been significantly improved to $37.9\%$ and $36.4\%$ on ADE20K 100-50 and ADE20K 100-10, respectively. In addition, the fairness has been promoted due to the mIoU improvement of minor groups. We also study the impact of network backbones and cluster margin $\nabla$ in our appendix.
Effectiveness of Scaling Factor of Cluster. Table 3 illustrates the experimental results of the impact of different scaling factor $\alpha$ on ADE20K 100-50 and Pascal VOC 15-5 benchmarks. As shown in Table 3, when the value of scaling factor $\alpha$ gradually decreases, the performance of our proposed approach is improved accordingly since the fairness contrastive loss in Eqn (4) tends to be more uniform across major and minor classes. However, when the scaling factor is too small ($\alpha=0.005$), the impact of the loss enforcement becomes weaker leading to the weaker enforcement of the fairness contrastive clustering, resulting in lower overall performance. In addition, we have observed that the higher the number of classes demands the higher the value of $\alpha$ since it will increase the compactness of more clusters.
Effectiveness of Loss Contributions. Table 4 illustrates the contributions of proposed learning objectives. For the model without using visual grammar, we only use a single unknown cluster ($N_{U}=1$) and adopt the nearest cluster strategies to assign clusters of unknown pixels. By using only cross-entropy loss, the mIoU performance remains low due to catastrophic forgetting and background shift problems. Meanwhile, with our fairness clustering loss $\mathcal{L}^{\alpha}_{Cont}$, visual grammar model $\phi$, and the cluster regularizer $\mathcal{R}$, the mIoU performance has been significantly improved to $37.9\%$ and $36.4\%$ on ADE20K 100-50 and ADE20K 100-10, respectively. Moreover, our FALCON has significantly promoted the fairness of segmentation models illustrated by the mIoU improvement of major and minor groups.
Effectiveness of Visual Grammar. We evaluate FALCON under three settings, i.e., Nearest Cluster, Fixed $\phi$ pretrained ImageNet (without updating on each learning step), and Adaptive $\phi$ (with updating on each learning step). As in Table 2, the mIoU result using only the nearest cluster remains ineffective. Meanwhile, the adaptive visual grammar model updated at each learning step further boosts the mIoU performance and promotes fairness, i.e., increased by $4.5\%$ and $4.9\%$ on ADE20K 100-50 and ADE20K 100-10 compared to the nearest cluster approach. In addition, we study the impact of choosing the number of features $M$ in the visual grammar model in our appendix. Fig. 6 illustrates the feature distributions of unknown classes (future class). As a result, our FALCON approach is able to model features of unknown classes into different clusters and produce better and more compact clusters compared to the one without Fairness Learning via Contrastive Attention.
Figure 6: Cluster Distribution at Learning Step $t=1$ of ADE20K 100-50 (classes 109-144 are future classes) without (left) and with (right) Fairness Learning via Contrastive Attention.
Effectiveness af Forgetting Each Learning Step. Table 5 reports the results of our model evaluated at each learning step. The performance of our model remains stable at each learning step. In particular, the maximum performance drop at learning step $4$ of classes 111-120 due forgetting is only $0.8\%$. Meanwhile, at the last learning task, the mIoU results of old classes are well maintained, with a slight drop of 0.4% in classes 131-140. As shown in the result, our model has effectively maintained the performance to avoid catastrophic forgetting while enabling the capability of learning classes.
Table 5: Effectiveness of CSS Model At Each Learning Step.
ADE20K 100-10
0-100 101-110 111-120 121-130 131-140 141-150 all
Step 1 44.9 -- -- -- -- -- 44.9
Step 2 44.8 15.8 -- -- -- -- 42.2
Step 3 44.7 15.7 22.1 -- -- -- 40.3
Step 4 44.6 15.6 21.3 26.3 -- -- 39.2
Step 5 44.5 15.4 21.2 26.1 24.4 -- 37.9
Step 6 44.4 15.4 21.2 26.1 24.0 15.5 36.4
avg 40.1
Table 6: Comparison with Prior Methods on ADE20K Benchmarks (Note: The results of MiB [3], PLOP [21], and FairCL [63] using Transformer on ADE20K 100-5 were not reported in prior studies. The upper bound results are not trained with fairness objective).
Network Method ADE20K 100-50 ADE20K 100-10 ADE20K 100-5
0-100 101-150 all avg 0-100 101-150 all avg 0-100 101-150 all avg
DeepLab-V3 PLOP [21] 41.9 14.9 32.9 37.4 40.5 14.1 31.6 36.6 39.1 7.8 28.8 35.3
RCIL [76] 42.3 18.8 34.5 -- 39.3 17.6 32.1 -- 38.5 11.5 29.6 --
REMINDER [46] 41.6 19.2 34.1 -- 39.0 21.3 33.1 -- 36.1 16.4 29.5 --
RCIL+LGKD [74] 43.3 25.1 37.2 -- 42.2 20.4 34.9 -- -- -- -- --
FairCL [63] 43.4 24.0 37.0 40.5 41.7 20.4 34.7 39.0 -- -- -- --
FALCON 44.6 24.5 37.9 41.3 44.4 20.4 36.4 40.1 38.0 16.1 30.7 37.6
[…]
Mask2Former MiB [3] 37.0 24.1 32.6 38.3 23.5 10.6 26.6 29.6 21.0 6.1 16.1 27.7
PLOP [21] 44.2 26.2 38.2 41.1 34.8 15.9 28.5 35.2 33.6 14.1 27.1 33.6
CoMFormer [4] 44.7 26.2 38.4 41.2 40.6 15.6 32.3 37.4 39.5 13.6 30.9 36.5
ECLIPSE [32] 45.0 21.7 37.1 -- 43.4 17.4 34.6 -- 43.3 16.3 34.2 --
CoMasTRe [26] 45.7 26.0 39.2 41.6 42.3 18.4 34.4 38.4 40.8 15.8 32.6 38.6
MiB [3] 43.4 30.6 39.2 38.7 39.1 20.4 34.2 39.5 -- -- -- --
PLOP [21] 43.8 26.2 38.0 38.1 43.3 24.1 36.2 40.3 -- -- -- --
Transformer FairCL [63] 43.6 25.5 37.6 40.7 42.2 21.9 35.5 39.4 -- -- -- --
FALCON 47.5 30.6 41.9 43.5 47.3 26.2 40.3 42.8 40.8 18.9 33.5 38.1
[…]
Figure 7: The Comparison of Qualitative Results on ADE20K Between Our Approach and RCIL [76].
Table 7: Comparisons with Prior Methods on Pascal VOC.
Method Pascal VOC 15-5 Pascal VOC 15-1 Pascal VOC 10-1
0-15 16-20 all 0-15 16-20 all 0-10 11-20 all
DeepLab-V3 MiB [3] 76.37 49.97 70.08 38.00 13.50 32.20 20.00 20.10 20.10
PLOP [21] 75.73 51.71 70.09 65.10 21.10 54.60 44.00 15.50 30.50
RCIL [76] -- -- -- 70.60 23.70 59.40 55.40 15.10 34.30
FairCL [63] -- -- -- 72.00 22.70 60.30 42.30 25.60 34.40
SSUL [5] 77.82 50.10 71.22 77.31 36.59 67.61 71.31 45.98 59.25
FALCON 79.35 54.77 73.50 78.34 42.57 69.83 73.94 49.73 62.41
[…]
Transformer PLOP [21] 72.51 48.37 66.76 64.59 37.23 58.08 48.53 33.71 41.47
SSUL [5] 79.91 56.83 74.41 79.91 40.56 70.54 74.06 51.85 63.48
FairCL [63] -- -- -- 73.50 22.80 61.50 57.10 14.20 36.60
FALCON 81.20 58.04 75.69 78.71 47.54 71.28 74.92 52.54 64.26
[…]
Table 8: Comparisons on Cityscapes.
Method 11-5 11-1 1-1
DeepLab-V3 LWF-MC [49] 58.90 56.92 31.24
ILT [39] 59.14 57.75 30.11
MİB [3] 61.51 60.02 42.15
PLOP [21] 63.51 62.05 45.24
RCIL [76] 64.30 63.00 48.90
FairCL [63] 66.96 66.61 49.22
FALCON 70.74 69.75 55.24
[…]
Trans. FairCL [63] 67.85 67.09 55.68
FALCON 71.33 70.14 58.79
[…]
4.3 Comparison with Prior SoTA Methods
ADE20K. Table 6 presents our experimental results compared to prior CSS methods. By using DeepLab-V3, our approach has achieved SoTA performance, i.e., the mIoU results of $37.9\%$ and $+36.4\%$ on ADE20K 100-50 and ADE20K 100-10 benchmarks, higher than prior FairCL [63]. Meanwhile, FALCON using Transformer has outperformed the prior CoMFormer [4] model by $+3.5\%$, $+8.0\%$, and $+2.6\%$ on ADE20K 100-50, ADE20K 100-10, and ADE20K 100-5. In addition, our mIoU results on the initial classes remain competitive with the upper-bounded results because FALCON can well handle the fairness problem compared to fully supervised learning. Although ECLIPSE [32] exhibits slightly better on ADE20K 100-5 by only $0.7\%$ (all) due to the better segmentation framework (i.e., Mask2Former), FALCON significantly outperforms ECLIPSE on ADE20K 100-50 and ADE20K 100-10. We also report our results of ADE20K 50-50 in the appendix. As in Fig. 7, FALCON produces better segmentation maps than prior methods.
Pascal VOC. Table 8 presents our results on Pascal VOC benchmarks. FALCON has consistently achieved the SoTA performance on three benchmarks. In particular, compared to the prior FairCL [63] approach, our methods using DeepLab-V3 have improved the mIoU performance up to $73.50\%$, $69.83\%$, and $62.41\%$ on Pascal VOC 15-5, Pascal VOC 15-1, and Pascal VOC 10-1, respectively. Additionally, by using the better backbone, the performance is further improved, and the gap with the upper-bound result is reduced.
Cityscapes. Table 8 reports the performance of our approach using DeepLab-V3 compared to prior methods on three different settings, i.e., Cityscapes 11-5, Cityscapes 11-1, and Cityscapes 1-1. As in Table 8, the performance of our methods has consistently outperformed prior FairCL [63] approach by $+3.78\%$, $+3.14\%$, and $+6.02\%$ on three benchmarks. Similar to our experiments on ADE20K and VOC, the better network brings higher results.
5 Conclusions and Limitations
Conclusions. This paper has presented a novel Fairness Learning via Contrastive Attention approach to CSS. In particular, the fairness contrastive clustering loss has been introduced to model both catastrophic forgetting and fairness problems. Then, the visual grammar model was presented to model the unknown classes. The experimental results on different benchmarks have shown the SoTA performance and fairness promotion of our proposed FALCON approach.
Limitations. Our study chose a set of learning hyper-parameters to support our theoretical analysis. However, it potentially consists of several limitations related to choosing learning parameters and cluster initialization. Further discussion of limitations is presented in the supplementary.
Acknowledgment This material is based upon work supported by the National Science Foundation under Award No. OIA-1946391. We also acknowledge the Arkansas High-Performance Computing Center for providing GPUs.